### v1

# BOP-D: Revisiting 6D Pose Estimation Benchmarks for Better Evaluation under Visual Ambiguities

Abstract
Currently, 6D pose estimation methods are benchmarked on datasets that consider, for their ground truth annotations, visual ambiguities as only related to global object symmetries. However, as previously observed [26], visual ambiguities can also happen depending on the viewpoint or the presence of occluding objects, when disambiguating parts become hidden. The visual ambiguities are therefore actually different across images. We thus first propose an automatic method to re-annotate those datasets with a 6D pose distribution specific to each image, taking into account the visibility of the object surface in the image to correctly determine the visual ambiguities. Given this improved ground truth, we re-evaluate the state-of-the-art methods and show this greatly modify the ranking of these methods. Our annotations also allow us to benchmark recent methods able to estimate a pose distribution on real images for the first time. We will make our annotations for the T-LESS dataset and our code publicly available.
Figure 1: We provide for the first time 6D pose annotations in the form of a per-image distribution. Current annotations in BOP [15] datasets are given as a single pose, shown here as a circle in the SO(3) representations. T-LESS [11] also provides a symmetry pattern leaving the 3D object model invariant, from which a distribution can be computed (the colored points in the SO(3) representations). Such distribution however does not cover many cases [26]: For the object in this example, when only its core is visible (Case 1), the pose is fully ambiguous and should be represented by a continuous distribution in SO(3). When the sides of the head are visible (Case 2), there are still ambiguities and the distribution is made of 6 clusters. When the hole is visible (Case 3), there are no ambiguities, and the pose distribution should be concentrated around one pose. Our method annotates the T-LESS dataset with per-image distributions, taking into account the partial occlusions and allowing us to evaluate a predicted pose properly. We show that considering these distributions for evaluating state-of-the-art methods results in a significant change of ranking for the BOP challenge.
1 Introduction
Visual 6D pose estimation of an object consists in determining the 3D position and orientation of this object with respect to the camera that explain the observed image. It is a key task in many application domains, such as robotics (e.g. grasping or manipulation), augmented reality, industrial quality control, etc. In such contexts, datasets of images with pose ground-truths are of primary importance, since they can be used to train learning based methods, but first and foremost to benchmark the performances of proposed approaches. The annotation accuracy of reference benchmarks, such as BOP [39] for 6D pose estimation, is thus crucial since it influences the research directions.
However, annotating images of objects with their 6D poses is a complex task. This is especially true when considered objects include symmetrical parts. Indeed, as illustrated in Figure 1, an image does not necessarily correspond to a single pose solution depending on the viewpoint and occlusions. While symmetrical objects naturally imply multiple solutions, non-symmetrical objects can also yield multiple solutions in case of partial occlusions. Current reference benchmarks [39] ignore these occlusion-induced symmetries, resulting in an imperfect evaluation of the methods. As we will show, considering better ground truth distributions results in a significant change of ranking of current pose estimation methods.
In this paper, we first propose a method that automatically estimates pose distributions for images that take into account the ambiguities due to the object symmetries but also to the partial occlusions. Our method exploits the current annotations as a starting point, the ground truth visibility masks of the objects, and their 3D models. While we demonstrate it on T-LESS, our method is fully automatic and can be applied to other datasets.
We also re-evaluate classical metrics with these annotations to evaluate the performance of pose estimation methods that return a single pose estimate per image. This yields a strongly modified ranking of current challenger methods.
Moreover, we also consider the few methods that already predict multi-modal pose distributions [7, 17]. Because of the absence of proper annotations, these methods have been evaluated quantitatively only on synthetic images and only qualitatively on real images. Thanks to our annotations, we are able to provide their first evaluation on real data.
In summary, our contributions are the following:
  1. 1.
    A novel automatic method computing a multi-modal 6D pose distributions ground truth from a unique ground truth pose and a proposal set of object symmetries;
  2. 2.
    A comprehensive re-evaluation of 6-DOF single pose estimation methods related to the re-annotation of the T-LESS [11] ground-truth;
  3. 3.
    A new evaluation framework of 6-DOF pose distribution estimation methods, which makes it the first multi-modal pose distribution benchmarking on real data.
Our annotations and code will be made public.
2 Related Work
Object 6D pose estimation aims at determining an object pose that best explains the camera observation. While the solution is unique when the object is non-symmetrical and not occluded, multiple equivalent possible solutions arise when the object is symmetrical or when the occlusions prevent the observation of symmetry-breaking elements.
While the state of the art mostly focused on estimating a single pose [4, 35, 13, 21, 38, 41, 6, 37], some recent works focus on estimating the complete set of solutions [26, 33, 30, 20, 27, 16, 7, 17].
As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are required, even for methods that returns a single pose per image.
Datasets and 6D pose annotation.
Various techniques have been introduced to annotate image datasets with their 6D poses. For synthetic datasets [2, 40], the ground truth pose is directly available, while, for real datasets, this pose is usually determined with the help of user interaction [5, 12, 3], markers [8, 9, 1], or a robotic arm [42, 28], and refined with ICP in the case of an RGB-D camera [5, 12, 42, 28].
However, such annotations only provide a single pose solution while multiple may exist. For fully symmetric objects, the solutions are always related with the same set of rigid transforms, independently of the viewpoint or the occlusions. This set of transforms, also called symmetries pattern, can then be pre-computed offline and applied to the initial solution to recover the whole set of solutions [14].
For non-symmetrical objects, recovering the complete set of solutions is much more challenging since symmetries may arise from the non-visibility of disambiguating parts of the object, those non-visibilities being induced by the viewpoint (self-occlusion), by other elements of the scene (external occlusion). The transformation set that relates the different solutions is then specific to each image and cannot be pre-computed from the 3D object model.
To our knowledge, no method has been proposed to determine this ground-truth transformation set per image. Instead, the current gold standard in 6D pose estimation benchmarks still consists in approximating this per-image transformation set with a unique global transformation set. The latter is usually computed with the same method than fully symmetric object with large enough tolerance to ignore small disambiguating elements (e.g. surface deviation tolerance of 1.5cm or 10% of the object diameter [14]).
Evaluation metrics.
Performance evaluation usually differs depending on whether the method is used to output a single pose or a distribution.
For single pose estimation, the accuracy is measured through the deviation of the object surface points when transformed by the estimate and by the ground truth pose (registration error) [10]. Depending on the sensor used—RGB or RGB-D—and the targeted application (e.g. robotics or Augmented Reality), deviations can be measured in 3D space (3D distance) or in the image space (reprojection error), and the registration error can be considered as the mean or maximal error over the object surface. In case of multi-valued ground-truths due to a symmetrical object or occlusion-induced symmetries, the accuracy corresponds to the minimal accuracy with respect to the set of solutions [10, 44, 40]. In practice, the most commonly used accuracy measures are :
  1. 1.
    Average Distance metric (ADD), and its variations for symmetrical objects (ADD-S [44], ADD-H [40]), target to measure the mean Euclidean error between the surface points of the object transformed with the estimated pose and the ground truth pose, but have been replaced by MSSD now.
  2. 2.
    Visual Surface Discrepancy (VSD) [12] measures the ratio of pixels for which the discrepancy between their value in the distance map computed with the ground truth pose and their values in the distance map computed with the pose to evaluate. VSD is more expensive to compute than MSSD and MSPD, and requires a depth image. It is now omitted for new tasks as stated by the BOP organizers on their website [^1] https://bop.felk.cvut.cz/method_info/447/.
  3. 3.
    Maximum Symmetry-Aware Surface Distance (MSSD) [14] which is similar to ADD but considers the maximal error instead of the mean error and symmetry management. MSSD provides a 3D error, useful to robotics applications.
  4. 4.
    Maximum Symmetry-Aware Projection Distance (MSPD) [14] measures the maximum reprojection error between the projection of the visible points with the ground-truth pose and their corresponding projection with the pose to evaluate and symmetry management. MSPD provides a reprojection error, useful to augmented reality applications.
Based on these accuracy measures, and inspired from the evaluation of the detection methods, single 6D pose estimation methods are usually evaluated through Precision/Recall. In such evaluation frameworks, an estimated pose is considered as correct if its registration error is below a predefined threshold. The precision corresponds to the rate of correct pose estimations, meaning the ratio of the estimated poses that are correct over the number of estimated poses. The recall corresponds to the rate of correctly registered objects, meaning the ratio between the number of object instances in the dataset whose pose was considered as correct over the number of object instances in the dataset.
Regarding the evaluation of multi-modal 6D pose distribution estimation methods, the accuracy is usually measured as a pose error, meaning the minimal deviation between the inferred rotation and translation and the corresponding nearest ground truth poses [30, 7, 18, 36].
If the method outputs a probability distribution over the whole pose space, the method is usually evaluated through its spread, corresponding to the expectation of the pose error and the log-likelihood between the inferred distribution and the multi-valued solution [30, 7, 18, 36, 16], meaning the sum of the mean log probability at ground truth solutions. While the spread provides a probabilistic measure of the accuracy, the log-likelihood measures how similar the probability distributions are. This penalizes the method if some ground truth solutions are missing in the estimate.
However, not all multi-modal 6D pose estimation methods actually provide a full probability distribution over the whole pose space. Instead, some methods output a set of poses corresponding to some local maxima of the underlying probability distribution [17]. For such methods, performances can be evaluated through the Precision/Recall of the multi-valuated estimation. A pose estimate is then considered as correct with respect to one ground truth pose if its distance in pose space do not exceed a threshold $\delta$. The Precision for a given image is then defined as the ratio of the number of estimated poses whose distance have at least one ground truth pose that do not exceed the threshold $\delta$ over the total number of estimated poses. The Recall for an object image is then defined as the ratio between the number of ground truth pose whose distance have at least one estimated pose that do not exceed the threshold $\delta$ over the total number of ground truth poses. Similarly to the spread and log likelihood, these precision and recall are related to the multi-modal pose distribution.
Our method combines a ground truth pose and a symmetries pattern to represent the pose distribution for each image. Unlike previous works, the symmetries pattern is adjusted to each image, taking into account its specific viewpoint and occluded objects (Section 3). Moreover, whereas single pose and multi-modal distribution pose methods are currently evaluated with non-comparable metrics, we introduce new evaluation metrics for multi-modal methods that are homogeneous with single pose methods (Section 4).
[…]
3 Method
As shown in Figure 2, the probability of a pose explaining a given image of the object is related to the visibility of the different points on the surface of the object.
More formally, given an image $I$ and a ground truth pose $P_{\text{GT}}$ for a target object in this image, the objective of our method is to recover the set ${\cal T}=\{T_{i}\in{\text{SE}}(3)\}$ of rigid transformations that transform the ground truth pose $P_{\text{GT}}$ into another pose that results in the same appearance of the object. In the following, we will call set ${\cal T}$ the symmetry pattern of image $I$.
Depending on the geometry $M$ of the object and the set ${\cal V}(M)$ of its visible points, set ${\cal T}$ can be limited to the identity transform in the case of a non=ambiguous pose or can include multiple transformations in addition to the identity. As introduced in [10], ${\cal T}$ can be defined as partial $\epsilon$-symmetries of ${\cal V}(M)$:
$\epsilon\text{-sym}({\cal V}(M),M)=\{T_{i}:d(T_{i}*{\cal V}(M),M)<\epsilon\}\>,$ (1)
where $d$ is the distance between the surfaces and $\epsilon$ the surface deviation tolerance.
If the set ${\cal V}(M)$ of visible points is available, the symmetry pattern could in principle be approximated by sampling the rigid transformation space SE(3) and assessing whether the sample belongs to the $\epsilon$-symmetries. However, since sampling a 6D space becomes quickly computationally expensive, such annotation process was never used at our knowledge.
In the following sections, we introduce solutions to reduce the computational cost (Section 3.1), how to introduce some tolerance in the evaluation of the $\epsilon$-symmetries (Section 3.2), and how to extend the annotation to textured object (Section 3.3).
3.1 Speeding up the computation of $\epsilon$-sym
Computing $\epsilon$-symmetries for each instance of image of a dataset can quickly become time consuming since it implies to sample the 6D pose space around the ground truth pose. To accelerate the computation of the $\epsilon$-symmetries, we reformulate them to introduce image-independent terms that can be pre-computed.
First, by definition of the $\epsilon$-symmetries, for any subsets ${\cal V}_{1}$ and ${\cal V}_{2}$ of visible points of a 3D model $M$, we have:
$\epsilon\text{-sym}({\cal V}_{1}\cup{\cal V}_{2},M)=\epsilon\text{-sym}({\cal V}_{1},M)\cap\epsilon\text{-sym}({\cal V}_{2},M)\>.$ (2)
Therefore, for any subset ${\cal V}=\{v_{j}\}$ of visible points of a 3D object $M$, we have:
$\epsilon\text{-sym}({\cal V},M)=\bigcap_{j}\epsilon\text{-sym}(v_{j},M)\>,$ (3)
where $\epsilon\text{-sym}(v_{j},M)$ is the symmetry pattern when the point $v_{j}$ is the only visible point, which we name elementary symmetry pattern of $v_{j}$.
With such expression of the $\epsilon$-symmetries, it appears that elementary symmetry patterns can be pre-computed for the whole 3D object model $M$ since they are independent of the image. Only the set of visible points and the intersection of the corresponding pre-computed elementary symmetries patterns need to be computed per image.
3.2 Robust symmetries patterns intersection
In practice, using a strict intersection might reject too many transformations since a transformation $T_{i}$ would be kept only if it is part of all the elementary symmetries patterns. Instead, we prefer to use a soft intersection, meaning that the transformation $T_{i}$ will be included even if a few patterns do not include $T_{i}$. This soft intersection is achieved by counting, for each transformation $T_{i}$, the number of elementary symmetries patterns that include $T_{i}$. This counting is defined as follows, in the form of a histogram over ${\cal T}$:
$\forall i,\text{H}(T_{i})=\text{Card}\{v_{j}~{}|~{}v_{j}\in{\cal V}(M),T_{i}\in\epsilon\text{-sym}(v_{j},M)\}\>.$ (4)
A strict intersection would only keep the $T_{i}$ with the maximum value of $\text{H}(T_{i})$ which is represented by $\text{H}(Id)$ as the identity of SE(3) and is always in the symmetry pattern. Our soft intersection tolerates a threshold and keeps all the $T_{i}$’s such that $\text{H}(T_{i})>\text{H}(Id)-\tau$, where $\tau$ represents the minimum size (expressed in number of 3D points) of the symmetry-breaking element.
3.3 Generalization to textured 3D objects
While the process previously described considered only texture-less 3D objects, the method can be extended to textured ones by simply redefining the $\epsilon$-symmetries as:
$\epsilon\text{-sym}({\cal V}(M),M)=\{T_{i}:d(T_{i}*{\cal V}(M),M)<\epsilon,\\ d_{\text{color}}(T_{i}*{\cal V}(M),M)<\zeta)\}\>,$ (5)
where $d_{\text{color}}$ is the distance in color-space and $\zeta$ is the color deviation tolerance.
4 Evaluating 6D Pose Distribution Predictions
Currently, single pose estimation methods and distribution pose estimation methods are evaluated on different datasets (BOP challenge for single poses, SYMSOL for 6D pose distributions) with different metrics (registration error for single pose estimation, pose error for 6D pose distribution).
With our annotation method that provides the full set of 6D pose solutions even on real data, it becomes possible to evaluate both tasks on the same dataset. However, it would be beneficial if both tasks can be evaluated with comparable metrics.
We therefore propose to keep the evaluation process of single pose method unchanged, but on a more accurate ground truth, and to extend it to pose distributions evaluation.
First, whereas error pose is commonly used in 6D pose distribution evaluation process [30, 18, 7], we propose to use the registration error. The latter takes the object 3D shape into account, making the accuracy more meaningful for most of the application. It is also the type of error used by the current gold standard benchmark for single pose [39]. Typically, we suggest to use the Maximal Surface Distance (MSD) and Maximum Projective Distance (MPD) metrics.
Second, we propose to keep the metrics of Precision and Recall, but we adapt them to measure Precision over the whole estimated distribution of poses instead of a unique pose, and Recall over the whole set of solutions instead of considering a pose to be found if one of its multiple solutions was found.
Precision for pose distribution.
We define the precision in the case of pose distribution prediction as:
$\mathbf{P_{d}}(\text{Est},\text{GT})=\\ \sum\limits_{P_{\text{Est}}^{j}\in Est}p(P_{\text{Est}}^{j})\left(\min_{P_{\text{GT}}\in\text{GT}}d(P_{\text{Est}}^{j},P_{\text{GT}})<\tau_{\mathbf{d}}\right)\>,$ (6)
where Est and GT are respectively the estimated and the ground truth distributions to compare, and $p(\text{Est}_{i})$ represents the probability associated to the $i$-th element of the estimated distribution  [^2] If the method output solutions without assigning a probability to each pose, a uniform probability over the solutions is used. , $\mathbf{d}$ is the chosen registration distance, and $\tau_{\mathbf{d}}$ the associated threshold.
Recall for pose distribution.
Similarly to the precision, we define the recall in the case of pose distribution prediction as:
$\mathbf{R_{d}}(\text{Est},\text{GT},\tau_{d})=\\ \sum\limits_{P_{\text{GT}}^{i}\in\text{GT}}\max\left(p(\hat{P}_{\text{Est}}^{i}),\frac{1}{\text{Card}(\text{GT})}\right)[d(\hat{P}_{\text{Est}}^{i},P_{\text{GT}}^{i})<\tau_{\mathbf{d}}],\\ \text{with }\hat{P}_{\text{Est}}^{i}=\operatorname*{arg\,min}_{P_{\text{Est}}\in\text{Est}}d(P_{\text{Est}},P_{\text{GT}}^{i})\>.$ (7)
The probability of $\hat{P}_{\text{Est}}^{i}$ is clamped to $\frac{1}{\text{Card}(\text{GT})}$ since ground truth poses are considered as equiprobable.
5 Experiments
In this section, after evaluating the performances of our annotation method (Section 5.2), we present and discuss the impact of our ground-truth annotations and evaluation metrics onto the evaluation of state-of-the-art solutions for both single pose estimation (Section 5.2) and multi-modal pose distribution estimation (Section 5.4).
Since our annotation process and metrics are related to objects with intrinsic or occlusion-induced symmetries, we perform our evaluations on the T-LESS dataset. Indeed, among the datasets of the BOP challenge [15], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground-truth annotations: ITODD [5] and Homebrewed [19] feature symmetrical objects, with occlusion-induced symmetries, but do not share publicly their ground-truths. HOPE [40], IC-BIN [3] and YCB-V [44] have symmetrical shapes, but texture information disambiguate them completely.
The first experiment focus on a qualitative evaluation of our annotation method (Section 5.2). The second on evaluate the impact of our re-annotation of T-LESS Dataset on the performance evaluation of the best performer methods of the BOP benchmark [15] on single pose estimation  (the Multiple instances of Multiple known objects (MiMo) task of BOP) (Section 5.3). The last experiment evaluate for the first time methods of 6D pose distribution estimation on a non-synthetic dataset with proper annotations and metrics (Section 5.4).
5.1 Implementation details
[…]
The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points.
5.2 Qualitative validation of the method
For the qualitative evaluation of our annotating method, we illustrate the results obtained on some images of the T-LESS dataset in Figure 3.
As expected, we can observe the coherence between the distribution and the visibility of the disambiguating elements. Since this dataset does not contain any image with visual ambiguity for the translational part of the pose, the ground truth translation is always reduced to a single 3D point.
More visualizations are provided in the supplementary material to convince the reader of the quality of our annotations.
Figure 3: Visualizations of our ground truth. We display SE(3) ground truth distribution for scenes 12, 16 and 17. Circle on orientation diagram represent the unique ground truth pose that is provided as input to our method. Colors are used to link object instances to their distributions. More visualizations are provided in the supplementary material to convince the reader of the quality of our annotations.
BOP [39] BOP-D (our annotations)
Methods Modality $\mathbf{MSSD}$ $\mathbf{MSPD}$ $\mathbf{Mean}$ Rank $\mathbf{MSSD}$ $\mathbf{MSPD}$ $\mathbf{Mean}$ Rank
T-LESS [11] Challenge: pose estimation
gpose2023 [45] RGB-D 92.1 94.6 93.4 1 86.5 89.1 87.8 1
gdrnppv2-rgbd-pbrreal [41, 25] RGB-D 90.2 92.6 91.4 2 84.8 87.2 86.0 3 ( $\downarrow$1)
gpose2023-pbr [41, 25] RGB-D 90.3 92.6 91.4 3 84.7 87.2 86.0 2 ( $\uparrow$1)
modalocclusion-rgbd RGB-D 89.7 92.9 91.3 4 83.2 86.4 84.8 4
gdrbpp v1.5 [41, 25] RGB-D 88.7 91.5 90.1 5 58.1 61.9 60.0 14 ( $\downarrow$9)
gdrnpp v1.4 [41, 25] RGB-D 88.4 90.9 89.6 6 60.6 63.7 62.2 10 ( $\downarrow$4)
gdrnpp v1.3 [41, 25] RGB-D 87.1 90.3 88.7 7 57.5 61.2 59.3 16 ( $\downarrow$9)
zebraposesat-effnetb4 [37] RGB-D 87.3 90.2 88.7 8 60.2 63.4 61.8 11 ( $\downarrow$3)
gdrnpp-pbr-rgbd-mmodel v1.2 [41, 25] RGB-D 87.2 90.1 88.7 9 60.8 64.2 62.5 9
hccepose RGB 81.4 94.8 88.1 10 53.1 64.5 58.8 18 ( $\downarrow$8)
hipose-cvpr24 [24] RGB-D 86.6 88.8 87.7 11 58.0 60.8 59.4 15 ( $\downarrow$4)
cosypose [21] RGB 83.6 90.7 87.2 12 57.2 64.0 60.6 13 ( $\downarrow$1)
pfa RGB-D 85.5 87.7 86.6 13 59.2 62.0 60.6 12 ( $\uparrow$1)
hccepose RGB 79.2 93.2 86.2 14 51.4 63.1 57.2 20 ( $\downarrow$6)
zebraposesa [37] RGB 79.1 93.0 86.1 15 51.8 63.4 57.6 19 ( $\downarrow$4)
gdrnppde v1.1 [41] RGB-D 78.4 92.2 85.3 16 69.9 82.9 76.4 7 ( $\uparrow$9)
leroy-fuseocclu-rgb RGB-D 78.0 91.7 84.8 17 70.4 83.3 76.9 6 ( $\uparrow$11)
gdpnpp v1.0 [41, 25] RGB-D 83.4 85.6 84.5 18 57.5 60.2 58.9 17 ( $\uparrow$1)
surfemb [6] RGB-D 82.9 85.9 84.4 19 75.8 78.7 77.2 5 ( $\uparrow$14)
poseio RGB-D 81.7 86.0 83.8 20 69.3 73.6 71.4 8 ( $\uparrow$12)
T-LESS [11] Best mode of pose distribution methods
SpyroPose [7] RGB 61.2 75.3 68.3 1 57.8 71.6 64.7 1
LiePose Diffusion [17] RGB 60.1 92.2 76.1 2 48.3 76.1 62.2 2
T-LESS [11] Challenge: pose estimation of unseen objects
foundationpose [43] RGB-D 62.9 64.1 63.5 1 56.7 58.0 57.4 1
foundpose [34] RGB 55.0 62.3 58.6 2 45.5 52.1 48.8 5 ( $\downarrow$3)
gigaposemegapose [32, 22] RGB 54.3 62.0 58.1 3 45.3 52.3 48.8 4 ( $\downarrow$1)
gigaposegenflowkabsch [32, 29] RGB-D 56.3 58.6 57.5 4 49.4 51.7 50.6 2 ( $\uparrow$2)
gigaposemegapose [32] RGB 51.5 59.2 55.3 5 38.5 45.1 41.8 11 ( $\downarrow$6)
genflow-multihypo [29] RGB 50.9 57.9 54.4 6 45.9 52.4 49.1 3 ( $\uparrow$3)
genflow-multihypo [29] RGB 50.4 57.3 53.9 7 44.9 51.4 48.1 6 ( $\uparrow$1)
genflow-multihypo16 [29] RGB-D 52.8 54.8 53.8 8 46.7 48.6 47.7 7 ( $\uparrow$1)
sam6d [23] RGB-D 52.9 54.5 53.7 9 39.7 41.2 40.4 15 ( $\downarrow$6)
sam6d [23] RGB-D 52.1 53.5 5208 10 38.3 39.7 39.0 18 ( $\downarrow$8)
genflow-multihypo [29] RGB-D 51.6 53.6 52.6 11 45.2 47.2 46.2 8 ( $\uparrow$3)
foundpose [43] RGB 49.1 55.6 52.4 12 37.7 43.4 40.6 13 ( $\downarrow$1)
foundpose [43] RGB 49.1 55.6 52.3 13 37.7 43.4 40.5 14 ( $\downarrow$1)
cnos-fastsam10 [31] RGB 48.5 55.9 52.2 14 40.6 47.1 43.8 10 ( $\uparrow$4)
Table 1: Results of BOP 2023 on T-LESS [11]. We report the top contenders results of the pose estimation challenge (2023 edition) evaluated with $\mathbf{MSSD}$ and $\mathbf{MSPD}$ on both the official ground truth and our ground truth. We rank the methods with the mean of $\mathbf{MSSD}$ and $\mathbf{MSPD}$ (similar to BOP, except we exclude $\mathbf{VSD}$ as BOP starts to abandon it) and re-rank them with the mean of $\mathbf{MSSD}$ and $\mathbf{MSPD}$ computed on our ground truth distributions. This produces drastic changes in the rankings. We also show the same scores for the best mode of distribution estimation methods, however, both of them use ground truth bounding boxes instead of a detector, so direct comparison would be unfair. Finally, we apply the same study to the pose estimation of unseen objects challenge. Similarly, the rankings change. This highlights the importance to consider image level symmetries as we proposed, to evaluate accurately pose estimation methods. (names of the methods are automatically processed from BOP csv result files)
5.3 Single Pose Estimation Evaluation
Since May 2023, raw results of BOP Challenge [15] public submissions have become publicly available to allow in-depth analysis. In this experiment, we use them to compare the performance evaluations of state-of-the-art methods on T-LESS for both the original 6D annotations and our annotations. We considered top contenders of the BOP challenge for the pose estimation and pose estimation for unseen objects tasks. We also included two methods for the multi-modal pose distribution estimation task [7, 17] for which we kept only the mode with highest confidence mode. We consider their full predicted distributions in the next subsection. Since these methods are not part of BOP challenge, we retrained them on the T-LESS train set including both photo-realistic synthetic images and real images.
Experimental protocol.
We evaluate the current (July 2024) top contenders of the BOP Challenge on the T-LESS dataset [11] on the pose estimation task and the pose estimation of unseen object task together with [7, 17].
This evaluation is based directly on the raw pose results provided by the contenders and the official evaluation scripts of BOP challenges, both being publicly available on the Challenge website. Following the BOP Challenge, the evaluation is based exclusively on recall. Our method ranking is based on the mean of Recalls computed on MSPD and MSSD accuracies. As discussed in Section 2, we do not consider VSD metric as it is now omitted for new tasks in BOP.
Results.
Table 1 reports the results of our evaluations. We can observe a large impact of the ground truth on the methods ranking since only the first place in the ranking remains unchanged. For example, regarding the ranking of the pose estimation task, the method Surfemb [6] moves from the 19th place with the original ground truth annotation to the 5th place with our annotation, gaining 14 places. For pose estimation of unseen objects, the difference of ranking can reach 8 places.
We observe large impacts on the recall performances as well. Whereas all the 20 first contenders of pose estimation task have a mean recall up to 80% and even up to 90% for 5 methods with the original annotations, only the 4 methods exceed this score with our annotation, with a majority of method having less than 65%. Similarly, all the 14 top contenders for the pose estimation of unseen objects task have a mean recall that exceed 50% with original annotation whereas only the best one exceed this score once evaluated with our annotations.
It is also interesting to note that the ranking order of variants 1.3, 1.4 and 1.5 of GDRNPP (based on different object detectors, training augmentations, and refinement methods) is changed when evaluated with our ground truth. It underlines that inaccurate ground truth annotation can induce erroneous conclusions in ablation studies.
5.4 Pose Distribution Evaluation
In this experiment, we provide the first evaluation of 6D pose distribution estimation methods on a dataset of real images.
Experimental protocol
Similarly to Section 5.3, the methods were trained on the train set of T-LESS that includes both photo-realistic synthetic images and real images.
The ground truth used for this dataset is the same than the one of Section 5.3 and is constituted of a discreet set of 6D pose obtained with the 6D pose annotation process introduced in Section 3. The performances of the methods are evaluated in Precision and Recall, using accuracy measures MPD and MSD as introduced in Section 4.
More details on methods processing and results formatting are given in the supplementary material.
Results
We evaluated SpyroPose [7] and LiePose-Diffusion [17] since they provide 6D pose distribution, unlike [30, 16] that provide distribution only over the rotation.
Quantative results are reported on Table 2 whereas qualitative results are provided in the supplementary material.
It appears that LiePose-Diffusion outperforms SpyroPose. Without surprise, performances decrease when MSD accuracy is used since those two methods use RGB images only.
Methods $\mathbf{P_{MPD}}$ $\mathbf{R_{MPD}}$ $\mathbf{P_{MSD}}$ $\mathbf{R_{MSD}}$
SpyroPose [7] 69.9 69.9 55.2 65.9
LiePose-Diff [17] 75.4 94.1 48.1 85.3
Table 2: Comparisons of pose distribution estimation methods on T-LESS using our ground truth pose distributions. So far, these methods were evaluated quantitatively on synthetic images only and only qualitatively on real images.
6 Limitations
While our solution improves the pose annotations, the resulting pose distribution depends on some parameters (tolerance factors, sampling of the SE(3) space, etc.). We selected their values carefully. We let the reader check the visualizations in the supplementary material.
Our annotation method relies on a geometric analysis of the scene. It is possible that the sensor resolution, sensor noise, field of view, or motion blur may also affect the visibility of the disambiguating parts. Considering these effects would improve even more the ground truth.
7 Conclusion
For simplicity, most of the 6D pose estimation benchmarks rely on a single 6D pose annotation per image, the latter being completed with a per-object symmetry pattern to transform this unique pose into a distribution. We argued that ignoring the per-image nature of the symmetry pattern is prone to bias in the resulting distribution ground truth and in the resulting performances evaluations.
We then proposed a method to annotate 6D pose distribution with a per-image analysis of the object symmetries. We illustrated that the resulting ground truth is more accurate. Moreover, when using this ground truth to re-evaluate current state-of-the-art methods, we showed that the ranking of these methods changes drastically. We also introduced metrics to evaluate methods that estimate a pose distribution and provided the first evaluation of such methods on real data.
We believe that these results should encourage the 3D vision community to adopt a new outlook on these methods and how well they work in practice, and should guide future research. To facilitate development, we will make our annotations and code publicly available.
Supplementary Material
1 Analysis of our BOP-D Annotations
1.1 Additionnal BOP-D Ground Truths Visualizations
We first provide more visualizations for qualitative appreciation of the new ground truth annotations accuracy in Figure 5.
These images are taken from a video compilation of all ground truth annotations sorted by object identifier, also provided as supplementary material (BOP_D_id239_supp_newGT_visualizations.mp4), to convince the reader of their quality. We invite the reader to stop on some frames and check that the distribution recovered by our method does correspond to the ambiguities in the image for the object in the bounding box.
1.2 Differences between Original Annotations and Ours
Figure 4 visualizes the differences between the poses that are accepted by BOP and the ones accepted when using our annotations. For this purpose, for each object, we provide a bar plot. This bar plot shows, for each image where the object appears, the percentage of poses considered correct by BOP that are also considered correct with our annotations. The bar plots show the percentages after sorting them, i.e., the bar on the left corresponds to the image with the smallest difference.
The annotations for some objects remain mostly unchanged. But because our method analyses more finely the possible object symmetries, many poses are actually not accepted with our annotations for the other objects.
Figure 4: Visualization of annotation changes compared to the T-LESS original annotations. Top: T-LESS objects with their identifier. Bottom: For each object, we plot the percentages of poses kept from the original annotations by our method over the images (sorted by percentages). T-LESS assumes full rotational symmetries, while our annotation method captures more complex symmetry patterns. Only Object 18 is perfectly symmetrical and our method retrieves the same poses as the original annotations. For the other objects, in particular the objects with complex symmetry patterns like the fist 4 objects. Our annotations significantly change the original annotations. The objects annotated as ’circular’ in BOP are highlighted in orange.
Figure 5: Visualizing some BOP-D ground truths as computed by our method. Each example features an object of interest, in the bounding box in the left image, and its BOP-D pose distribution, split between the rotation part (center) and translation part (right). As no object present symmetries in translation here, the translation part of the distribution is always on the same point of $\mathbb{R}^{3}$. We provide much more examples in the accompanying video.
2 Computing the Pose Estimation Results
2.1 T-LESS Pose Estimation and Unseen Objects Pose Estimation
Recently, the BOP challenge [39] made public the pose estimations results of the methods evaluated in the leaderboard  [^3] https://bop.felk.cvut.cz/leaderboards/. Based on these results and the BOP toolkit  [^4] https://github.com/thodan/bop_toolkit/tree/master in which we implemented the variations of MSSD and MSPD, that use our per-image symmetries patterns instead of BOP global object symmetries, we were able to reprocess these pose estimates against our new ground truths. Our results have been presented in Table 1. Section 3 of supplementary material illustrates some of the failure cases with the new and more accurate ground truth.
2.2 Computation of the SpyroPose [7] Distribution Results
The implementation provided by the authors of SpyroPose allows to train a network for one object of T-LESS. We batched the training stage for all objects, and then batched the inference. For one object, SpyroPose produces more than 100 000 estimates, sorted by their probabilities, as it samples SO(3) $\times$ $\mathbb{R}^{3}$ (using a slice of $\mathbb{R}^{3}$). As the probabilities of these estimates quickly tend to zero, we reduce their distributions to the 400 best estimates. These 400 estimates are used for the evaluations of Section 5.4.
As no pose distribution evaluation existed, the authors of [7] reported only per-object averaged log-likelihood against BOP original ground truth.
2.3 Computation of the Lie-Pose Diffusion [17] Distribution Results
The implementation provided by the authors of Lie-Pose allows to train a network for all objects of T-LESS. The inference phase produces one pose estimate per image crop. We batched the inference phase, with varying seeds. We then merged all these estimates into a single set of poses per image crop. The authors report 1000 runs to produce a distribution. For our evaluation, we ran the code 100 times with different noises to produce Lie-Pose distribution results.
As no pose distribution evaluation existed, the authors of [17] reported only few illustrations of pose distributions results on T-LESS. They evaluated their method as a single pose estimation method, with a single run of the method.
3 Illustrations of Single Pose Re-evaluation
We provide here more cues about the changes in the Single Pose ranking, based on the MSSD and MSPD metrics using our new more accurate ground truth.
To do so, we look at the pose estimates of the method GDRNPP v1.4. When we evaluated GDRNPP v1.4 estimates against our new ground truths, it changed the method ranking from rank 6 to rank 10 in Table 1. Even more interestingly, the metrics went down from an MSSD of 88.4 and an MSPD of 90.9 to an MSSD of 60.6 and an MSPD of 63.7.
Figure 6 illustrates why MSSD and MSPD changed for this method. It appears that, although GDRNPP v1.4 estimates were close to the ground truth pose, its rotations were not precise. When they are evaluated against a symmetries pattern that is not precise, the evaluation appears correct. Our new ground truth shows that GDRNPP v1.4 tends not to align correctly some of the objects. Hence the drop in performances.
Refer to caption Refer to caption
(a) BOP continuous symmetry pattern of Object 2 (envelop) and GDRNPP v1.4 estimate (plain circle). The circle belongs to the envelop, yielding a low MSSD error of 4.46. (b) Our visual symmetry pattern (the much smaller envelop) and GDRNPP estimate (plain circle). The circle does not belong to the envelop anymore, the MSSD error becomes 21.62.
(c) Corresponding image for (a) and (b): T-LESS Scene 1, Image 17, Object 2 (in bounding box).
Refer to caption Refer to caption
(d) BOP continuous symmetry pattern for Object 1 (envelop) and GDRNPP v1.4 estimate (plain circle). The circle belongs to the envelop, yielding a low MSSD error of 9.43. (e) Our visual symmetry pattern (much smaller envelop) and GDRNPP estimate (plain circle). The circle does not belong to the envelop anymore, the MSSD error becomes 32.87.
Refer to caption Refer to caption
BOP continuous symmetry pattern of Object 4 (envelop) and GDRNPP v1.4 estimate (plain circle). The circle belongs to the envelop, yielding a low MSSD error of 3.33. Our visual symmetry pattern (much smaller envelop) and GDRNPP estimate (plain circle). The circle does not belong to the envelop anymore, the MSSD error becomes 35.38.
(h) Corresponding image for T-LESS Scene 5, Image 70, Objects 1 (d-e) and and 4 (f-g), in bounding boxes).
Figure 6: Impact of our annotations on Single Pose evaluation. We show here cases where the estimates by a state-of-the-art method (GDRNPP v1.4) produces fairly good MSSD errors when considering ground truth provided by BOP. For these cases, our more accurate ground truth yields worse MSSD errors, as it appears that the estimate belongs to the global symmetry pattern, but does not explain what is visible in the image.
4 Visualizing results by SpyroPose [7]
We display some of SpyroPose distribution estimates against our ground truth in Figure 7. For the case of the three instances of Object 1, SpyroPose correctly retrieves the single mode for Instance 1. The continuous symmetry of Instance 2 is partially retrieved.
Figure 7: Illustration of SpyroPose [7] results.. We show the distribution estimates for the 3 instances of Object 1 (in the bounding box). The ground truth distribution is displayed as an envelop for the rotation part and as red star for the translation part.
5 Visualizing results by LiePose diffusion [17]
We show some of LiePose [17] distribution estimates against our ground truth in Figure 8. Similarly to SpyroPose [7], LiePose [17] is able to retrieve the single mode of Instance 1, but gets better results when estimating continuous distributions.
Figures 9, 10, 11 and 12 compare SpyroPose [7] and LiePose [17] results on objects with discrete and continuous symmetries. SpyroPose [7] rotations tends to be more precise than LiePose [17], but misses some of the modes. LiePose [17] tends to estimate continuous symmetries when the image produces discrete ones. These images are taken from a video compilation of SpyroPose [7] and LiePose [17] results also provided as supplementary material (BOP_D_id239_supp_distribution_comparison_SpyroPose_LiePose.mp4). Scenes with single instance of objects have been chosen, to facilitate visualization. We invite the reader to stop on some frames and check the differences in the estimates. Our ground truth distribution is displayed as the envelop for the rotation part and as red stars for the translation part.
Figure 8: Illustration of LiePose [17] results.. We show the distribution estimates for the 3 instances of Object 1 (in the bounding box). The ground truth distribution is displayed as an envelop for the rotation part and as red star for the translation part.
Figure 9: Visualizing SpyroPose [7] (top row) and LiePose [17] (bottom row) distribution results for object 27 (four rotation modes). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). Both methods are able to retrieve the four rotation modes of the object. The envelop in the rotation part represents our BOP-D annotation. We provide much more examples in the accompanying video.
Figure 10: Visualizing SpyroPose [7] (top row) and LiePose [17] (bottom row) distribution results for object 15 (continuous rotation). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). Both methods fail to generate the target continuous rotation, although SpyroPose produces more correct rotations. The envelop in the rotation part represents our BOP-D annotation. We provide much more examples in the accompanying video.
Figure 11: Visualizing SpyroPose [7] (top row) and LiePose [17] (bottom row) distribution results for object 1 (six rotation modes). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). For this case of six rotations modes, SpyroPose is able to retrieve only two of them, whereas LiePose tends to a continuous distribution, thus generating false rotations. The envelop in the rotation part represents our BOP-D annotation. We provide much more examples in the accompanying video.
Figure 12: Visualizing SpyroPose [7] (top row) and LiePose [17] (bottom row) distribution results for object 3 (eight rotation modes). Each example features an object of interest, in the bounding box in the left image, and the methods distribution estimation, split between the rotation part (center) and translation part (right). For this case of eight rotations modes, SpyroPose is able to retrieve only two of them, whereas LiePose tends to a continuous distribution, thus generating false rotations. The envelop in the rotation part represents our BOP-D annotation. We provide much more examples in the accompanying video.
6 Using Our Per-Image Annotations for Other Pose Estimation Datasets
Among BOP datasets, ITODD [5] and HomeBrewedDB [19] are the two other presenting object symmetries. They could easily be processed by our method, however their ground truth poses, needed as input to our method, are not public.
We give here an illustration of the interest to reprocess their symmetries patterns. We take ITODD’s small validation set, for which the ground truth is public. Figure 13 presents our result for the star object new ground truth, as well as one case of the current best performer (gpose2023) failing to align the holes. In the current version of BOP evaluation, this pose estimation is validated. With our annotations, it would be penalized.
Figure 13: Illustration of ITODD symmetries on the validation set (with public GT). For the star image (left, first row), BOP symmetries display 12 rotation modes (middle), whereas our method would keep only 2 rotation modes, which align the two holes (right) (size was set to one over the number of modes, hence the bigger modes on the right). We also show a pose estimate of GPose [45], ranked first at BOP 2023, for the star object overlayed on its image. We observe that the holes are not correctly aligned. MSPD and MSSD metrics validate this estimate, whereas MSPD and MSSD metrics with our new annotations would have penalized it.
Acknowledgment
The authors thank Rasmus Laurvig Haugaard for providing SpyroPose [7] implementation for baseline evaluation.

### v2

@@
-BOP-D: Revisiting 6D Pose Estimation Benchmarks for Better Evaluation under Visual Ambiguities
+BOP-Distrib: Revisiting 6D Pose Estimation Benchmarks for Better Evaluation under Visual Ambiguities
 Abstract
-Currently, 6D pose estimation methods are benchmarked on datasets that consider, for their ground truth annotations, visual ambiguities as only related to global object symmetries.
-However, as previously observed [26], visual ambiguities can also happen depending on the viewpoint or the presence of occluding objects, when disambiguating parts become hidden.
-The visual ambiguities are therefore actually different across images.
+6D pose estimation aims at determining the pose of the object that best explains the camera observation.
+The unique solution for a non-symmetrical object can turn into a multi-modal pose distribution for a symmetrical object or when occlusions of symmetry-breaking elements happen, depending on the viewpoint.
+Currently, 6D pose estimation methods are benchmarked on datasets that consider, for their ground truth annotations, visual ambiguities as only related to global object symmetries, whereas they should be defined per-image to account for the camera viewpoint.
 We thus first propose an automatic method to re-annotate those datasets with a 6D pose distribution specific to each image, taking into account the visibility of the object surface in the image to correctly determine the visual ambiguities.
-Given this improved ground truth, we re-evaluate the state-of-the-art methods and show this greatly modify the ranking of these methods.
-Our annotations also allow us to benchmark recent methods able to estimate a pose distribution on real images for the first time.
+Second, given this improved ground truth, we re-evaluate the state-of-the-art single pose methods and show that this greatly modifies the ranking of these methods.
+Third, as some recent works focus on estimating the complete set of solutions, we derive a precision/recall formulation to evaluate them against our image-wise distribution ground truth, making it the first benchmark for pose distribution methods on real images.
 We will make our annotations for the T-LESS dataset and our code publicly available.
-Figure 1: We provide for the first time 6D pose annotations in the form of a per-image distribution.
+Figure 1: We provide for the first time 6D pose annotations in the form of a per-image object pose distribution.
 Current annotations in BOP [17] datasets are given as a single pose, shown here as a circle in the SO(3) representations.
@@
 Such distribution however does not cover many cases [29]: For the object in this example, when only its core is visible (Case 1), the pose is fully ambiguous and should be represented by a continuous distribution in SO(3).
-When the sides of the head are visible (Case 2), there are still ambiguities and the distribution is made of 6 clusters.
+When the sides of the head are visible (Case 2), there are still ambiguities and the distribution is made of 6 modes.
 When the hole is visible (Case 3), there are no ambiguities, and the pose distribution should be concentrated around one pose.
@@
 We show that considering these distributions for evaluating state-of-the-art methods results in a significant change of ranking for the BOP challenge.
+Such ground truth distributions also become a key asset when it comes to evaluating pose distribution estimation methods [9, 19].
+With appropriate metrics, we demonstrate the first quantitative evaluation of pose distribution methods on real images, as an extension to single pose methods.
 1 Introduction
-Visual 6D pose estimation of an object consists in determining the 3D position and orientation of this object with respect to the camera that explain the observed image.
+Visual 6D pose estimation of an object consists in determining the 3D position and orientation of this object with respect to the camera that explains the observed image.
 It is a key task in many application domains, such as robotics (e.g.
 grasping or manipulation), augmented reality, industrial quality control, etc.
-In such contexts, datasets of images with pose ground-truths are of primary importance, since they can be used to train learning based methods, but first and foremost to benchmark the performances of proposed approaches.
+In such contexts, datasets of images with ground-truth poses are of primary importance, since they can be used to train learning based methods, but first and foremost to benchmark the performances of proposed approaches.
 The annotation accuracy of reference benchmarks, such as BOP [42] for 6D pose estimation, is thus crucial since it influences the research directions.
@@
 As we will show, considering better ground truth distributions results in a significant change of ranking of current pose estimation methods.
-In this paper, we first propose a method that automatically estimates pose distributions for images that take into account the ambiguities due to the object symmetries but also to the partial occlusions.
+In this paper, we first propose a method that automatically annotates pose distributions for images that take into account the ambiguities due to the object symmetries but also to the partial occlusions.
 Our method exploits the current annotations as a starting point, the ground truth visibility masks of the objects, and their 3D models.
-While we demonstrate it on T-LESS, our method is fully automatic and can be applied to other datasets.
-We also re-evaluate classical metrics with these annotations to evaluate the performance of pose estimation methods that return a single pose estimate per image.
+While we demonstrate it on T-LESS [13], our method is fully automatic and can be applied to other datasets as demonstrated on YCB-V [47].
+We also re-evaluate classical metrics with these annotations to assess the performance of pose estimation methods that return a single pose estimate per image.
 This yields a strongly modified ranking of current challenger methods.
@@
   1. 1.
-    A novel automatic method computing a multi-modal 6D pose distributions ground truth from a unique ground truth pose and a proposal set of object symmetries;
+    A novel automatic method computing a multi-modal 6D pose distribution ground truth from a unique ground truth pose and a proposal set of object symmetries;
   2. 2.
@@
 While the state of the art mostly focused on estimating a single pose [6, 38, 15, 23, 41, 44, 8, 40], some recent works focus on estimating the complete set of solutions [29, 36, 33, 22, 30, 18, 9, 19].
-As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are required, even for methods that returns a single pose per image.
+As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are required, even for methods that return a single pose per image.
 Datasets and 6D pose annotation.
@@
 robotics or Augmented Reality), deviations can be measured in 3D space (3D distance) or in the image space (reprojection error), and the registration error can be considered as the mean or maximal error over the object surface.
-In case of multi-valued ground-truths due to a symmetrical object or occlusion-induced symmetries, the accuracy corresponds to the minimal accuracy with respect to the set of solutions [10, 44, 40].
+In case of multi-valued ground-truths due to a symmetrical object or occlusion-induced symmetries, the accuracy corresponds to the minimal error with respect to the set of solutions [12, 47, 43].
 In practice, the most commonly used accuracy measures are :
   1. 1.
-    Average Distance metric (ADD), and its variations for symmetrical objects (ADD-S [44], ADD-H [40]), target to measure the mean Euclidean error between the surface points of the object transformed with the estimated pose and the ground truth pose, but have been replaced by MSSD now.
+    Average Distance metric (ADD), and variations for symmetrical objects (ADD-S [47], ADD-H [43]), target to measure mean Euclidean error between surface points of the object transformed with the estimated pose and the ground truth pose, but are replaced by MSSD now.
   2. 2.
-    Visual Surface Discrepancy (VSD) [12] measures the ratio of pixels for which the discrepancy between their value in the distance map computed with the ground truth pose and their values in the distance map computed with the pose to evaluate.
+    Visual Surface Discrepancy (VSD) [14] measures misalignment over the visible surface of the object model.
 VSD is more expensive to compute than MSSD and MSPD, and requires a depth image.
-It is now omitted for new tasks as stated by the BOP organizers on their website [^1] https://bop.felk.cvut.cz/method_info/447/.
+It is now omitted for new tasks as stated by BOP organizers [^1] https://bop.felk.cvut.cz/challenges/.
   3. 3.
@@
 The recall corresponds to the rate of correctly registered objects, meaning the ratio between the number of object instances in the dataset whose pose was considered as correct over the number of object instances in the dataset.
-Regarding the evaluation of multi-modal 6D pose distribution estimation methods, the accuracy is usually measured as a pose error, meaning the minimal deviation between the inferred rotation and translation and the corresponding nearest ground truth poses [30, 7, 18, 36].
+Regarding the evaluation of multi-modal 6D pose distribution estimation methods, the accuracy is usually measured as a pose error, meaning the minimal deviation between the inferred rotation and translation and the corresponding nearest ground truth pose [33, 9, 20, 39].
 If the method outputs a probability distribution over the whole pose space, the method is usually evaluated through its spread, corresponding to the expectation of the pose error and the log-likelihood between the inferred distribution and the multi-valued solution [33, 9, 20, 39, 18], meaning the sum of the mean log probability at ground truth solutions.
@@
 For such methods, performances can be evaluated through the Precision/Recall of the multi-valuated estimation.
-A pose estimate is then considered as correct with respect to one ground truth pose if its distance in pose space do not exceed a threshold $\delta$.
-The Precision for a given image is then defined as the ratio of the number of estimated poses whose distance have at least one ground truth pose that do not exceed the threshold $\delta$ over the total number of estimated poses.
-The Recall for an object image is then defined as the ratio between the number of ground truth pose whose distance have at least one estimated pose that do not exceed the threshold $\delta$ over the total number of ground truth poses.
+A pose estimate is then considered as correct with respect to one ground truth pose if its distance in pose space does not exceed a threshold $\delta$.
+The Precision for a given image is then defined as the ratio of the number of estimated poses whose distance have at least one ground truth pose that does not exceed the threshold $\delta$ over the total number of estimated poses.
+The Recall for an object image is then defined as the ratio between the number of ground truth poses whose distance have at least one estimated pose that does not exceed the threshold $\delta$ over the total number of ground truth poses.
 Similarly to the spread and log likelihood, these precision and recall are related to the multi-modal pose distribution.
@@
 In the following, we will call set ${\cal T}$ the symmetry pattern of image $I$.
-Depending on the geometry $M$ of the object and the set ${\cal V}(M)$ of its visible points, set ${\cal T}$ can be limited to the identity transform in the case of a non=ambiguous pose or can include multiple transformations in addition to the identity.
+Depending on the geometry $M$ of the object and the set ${\cal V}(M)$ of its visible points, set ${\cal T}$ can be limited to the identity transform in the case of a non-ambiguous pose or can include multiple transformations in addition to the identity.
 As introduced in [12], ${\cal T}$ can be defined as partial $\epsilon$-symmetries of ${\cal V}(M)$:
 $\epsilon\text{-sym}({\cal V}(M),M)=\{T_{i}:d(T_{i}*{\cal V}(M),M)<\epsilon\}\>,$ (1)
-where $d$ is the distance between the surfaces and $\epsilon$ the surface deviation tolerance.
+where $d$ measures the mean deviation between the points of ${\cal V}(M)$ and $M$, $\epsilon$ is the maximal surface deviation tolerance and * represents the application of the transformation $T_{i}$ onto the point cloud ${\cal V}(M)$.
 If the set ${\cal V}(M)$ of visible points is available, the symmetry pattern could in principle be approximated by sampling the rigid transformation space SE(3) and assessing whether the sample belongs to the $\epsilon$-symmetries.
-However, since sampling a 6D space becomes quickly computationally expensive, such annotation process was never used at our knowledge.
-In the following sections, we introduce solutions to reduce the computational cost (Section 3.1), how to introduce some tolerance in the evaluation of the $\epsilon$-symmetries (Section 3.2), and how to extend the annotation to textured object (Section 3.3).
+However, since sampling a 6D space becomes quickly computationally expensive, such annotation process was never used to our knowledge.
+In the following sections, we introduce solutions to reduce the computational cost (Section 3.1), how to introduce some tolerance in the evaluation of the $\epsilon$-symmetries (Section 3.2), and how to extend the annotation to textured objects (Section 3.3).
 3.1 Speeding up the computation of $\epsilon$-sym
@@
 To accelerate the computation of the $\epsilon$-symmetries, we reformulate them to introduce image-independent terms that can be pre-computed.
-First, by definition of the $\epsilon$-symmetries, for any subsets ${\cal V}_{1}$ and ${\cal V}_{2}$ of visible points of a 3D model $M$, we have:
-$\epsilon\text{-sym}({\cal V}_{1}\cup{\cal V}_{2},M)=\epsilon\text{-sym}({\cal V}_{1},M)\cap\epsilon\text{-sym}({\cal V}_{2},M)\>.$ (2)
-Therefore, for any subset ${\cal V}=\{v_{j}\}$ of visible points of a 3D object $M$, we have:
-$\epsilon\text{-sym}({\cal V},M)=\bigcap_{j}\epsilon\text{-sym}(v_{j},M)\>,$ (3)
-where $\epsilon\text{-sym}(v_{j},M)$ is the symmetry pattern when the point $v_{j}$ is the only visible point, which we name elementary symmetry pattern of $v_{j}$.
+First, by definition of the $\epsilon$-symmetries, for any subsets $V_{1}\subset{\cal V}(M)$ and $V_{2}\subset{\cal V}(M)$, we have:
+$\epsilon\text{-sym}(V_{1}\cup V_{2},M)=\epsilon\text{-sym}(V_{1},M)\cap\epsilon\text{-sym}(V_{2},M)\>.$ (2)
+Therefore, we can reformulate Equation 1 as:
+$\epsilon\text{-sym}({\cal V}(M),M)=\bigcap_{v_{j}\in{\cal V}(M)}\epsilon\text{-sym}(v_{j},M)\>,$ (3)
+where $v_{j}$ is a 3D point and $\epsilon\text{-sym}(v_{j},M)$ is the symmetry pattern when the point $v_{j}$ is the only visible point, which we name elementary symmetry pattern of $v_{j}$.
 With such expression of the $\epsilon$-symmetries, it appears that elementary symmetry patterns can be pre-computed for the whole 3D object model $M$ since they are independent of the image.
@@
 A strict intersection would only keep the $T_{i}$ with the maximum value of $\text{H}(T_{i})$ which is represented by $\text{H}(Id)$ as the identity of SE(3) and is always in the symmetry pattern.
-Our soft intersection tolerates a threshold and keeps all the $T_{i}$’s such that $\text{H}(T_{i})>\text{H}(Id)-\tau$, where $\tau$ represents the minimum size (expressed in number of 3D points) of the symmetry-breaking element.
+Our soft intersection tolerates a threshold and keeps all the $T_{i}$ such that $\text{H}(T_{i})>\text{H}(Id)-\tau$, where $\tau$ represents the minimum size (expressed in number of 3D points) of the symmetry-breaking elements.
 3.3 Generalization to textured 3D objects
@@
 We therefore propose to keep the evaluation process of single pose method unchanged, but on a more accurate ground truth, and to extend it to pose distributions evaluation.
-First, whereas error pose is commonly used in 6D pose distribution evaluation process [30, 18, 7], we propose to use the registration error.
-The latter takes the object 3D shape into account, making the accuracy more meaningful for most of the application.
+First, whereas pose error is commonly used in 6D pose distribution evaluation process [33, 20, 9], we propose to use the registration error.
+The latter takes the object 3D shape into account, making the accuracy more meaningful for most of the applications.
 It is also the type of error used by the current gold standard benchmark for single pose [42].
@@
 $\mathbf{P_{d}}(\text{Est},\text{GT})=\\ \sum\limits_{P_{\text{Est}}^{j}\in Est}p(P_{\text{Est}}^{j})\left(\min_{P_{\text{GT}}\in\text{GT}}d(P_{\text{Est}}^{j},P_{\text{GT}})<\tau_{\mathbf{d}}\right)\>,$ (6)
-where Est and GT are respectively the estimated and the ground truth distributions to compare, and $p(\text{Est}_{i})$ represents the probability associated to the $i$-th element of the estimated distribution  [^2] If the method output solutions without assigning a probability to each pose, a uniform probability over the solutions is used.
-, $\mathbf{d}$ is the chosen registration distance, and $\tau_{\mathbf{d}}$ the associated threshold.
+where Est and GT are respectively the estimated and the ground truth distributions to compare, and $p(\text{Est}_{i})$ represents the probability associated to the $i$-th element of the estimated distribution  [^2] If the method outputs solutions without assigning a probability to each pose, such as [19], a uniform probability over the solutions is used., $\mathbf{d}$ is the chosen registration distance, and $\tau_{\mathbf{d}}$ the associated threshold.
 Recall for pose distribution.
@@
 Since our annotation process and metrics are related to objects with intrinsic or occlusion-induced symmetries, we perform our evaluations on the T-LESS dataset.
-Indeed, among the datasets of the BOP challenge [15], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground-truth annotations: ITODD [5] and Homebrewed [19] feature symmetrical objects, with occlusion-induced symmetries, but do not share publicly their ground-truths.
-HOPE [40], IC-BIN [3] and YCB-V [44] have symmetrical shapes, but texture information disambiguate them completely.
+Indeed, among the datasets of the BOP challenge [17], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground-truth annotations: ITODD [7] and Homebrewed [21] feature symmetrical objects, with occlusion-induced symmetries, but do not publicly share their ground-truths.
+HOPE [43] and IC-BIN [5] have symmetrical shapes, but texture information disambiguates them completely.
 The first experiment focus on a qualitative evaluation of our annotation method (Section 5.2).
-The second on evaluate the impact of our re-annotation of T-LESS Dataset on the performance evaluation of the best performer methods of the BOP benchmark [15] on single pose estimation  (the Multiple instances of Multiple known objects (MiMo) task of BOP) (Section 5.3).
-The last experiment evaluate for the first time methods of 6D pose distribution estimation on a non-synthetic dataset with proper annotations and metrics (Section 5.4).
+The second evaluates the impact of our re-annotation of T-LESS Dataset on the performance evaluation of the best performer methods of the BOP benchmark [17] on single pose estimation  (the Multiple instances of Multiple known objects (MiMo) task of BOP) (Section 5.3).
+The last experiment evaluates for the first time methods of 6D pose distribution estimation on a non-synthetic dataset with proper annotations and metrics (Section 5.4).
 5.1 Implementation details
@@
 The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object.
-For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points.
+For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 2.5$\times$2.5mm2.
 5.2 Qualitative validation of the method
-For the qualitative evaluation of our annotating method, we illustrate the results obtained on some images of the T-LESS dataset in Figure 3.
-As expected, we can observe the coherence between the distribution and the visibility of the disambiguating elements.
+We illustrate qualitatively the results obtained on some images of the T-LESS dataset in Figure 3.
+We can observe the coherence between the distribution and the visibility of the disambiguating elements.
 Since this dataset does not contain any image with visual ambiguity for the translational part of the pose, the ground truth translation is always reduced to a single 3D point.
@@
 Figure 3: Visualizations of our ground truth.
-We display SE(3) ground truth distribution for scenes 12, 16 and 17.
-Circle on orientation diagram represent the unique ground truth pose that is provided as input to our method.
+We display SE(3) ground truth distributions for scene 16 of T-LESS [13].
+Circle on orientation diagram represents the unique ground truth pose provided as input to our method.
 Colors are used to link object instances to their distributions.
 More visualizations are provided in the supplementary material to convince the reader of the quality of our annotations.
-BOP [39] BOP-D (our annotations)
-Methods Modality $\mathbf{MSSD}$ $\mathbf{MSPD}$ $\mathbf{Mean}$ Rank $\mathbf{MSSD}$ $\mathbf{MSPD}$ $\mathbf{Mean}$ Rank
-T-LESS [11] Challenge: pose estimation
-gpose2023 [45] RGB-D 92.1 94.6 93.4 1 86.5 89.1 87.8 1
-gdrnppv2-rgbd-pbrreal [41, 25] RGB-D 90.2 92.6 91.4 2 84.8 87.2 86.0 3 ( $\downarrow$1)
-gpose2023-pbr [41, 25] RGB-D 90.3 92.6 91.4 3 84.7 87.2 86.0 2 ( $\uparrow$1)
-modalocclusion-rgbd RGB-D 89.7 92.9 91.3 4 83.2 86.4 84.8 4
-gdrbpp v1.5 [41, 25] RGB-D 88.7 91.5 90.1 5 58.1 61.9 60.0 14 ( $\downarrow$9)
-gdrnpp v1.4 [41, 25] RGB-D 88.4 90.9 89.6 6 60.6 63.7 62.2 10 ( $\downarrow$4)
-gdrnpp v1.3 [41, 25] RGB-D 87.1 90.3 88.7 7 57.5 61.2 59.3 16 ( $\downarrow$9)
-zebraposesat-effnetb4 [37] RGB-D 87.3 90.2 88.7 8 60.2 63.4 61.8 11 ( $\downarrow$3)
-gdrnpp-pbr-rgbd-mmodel v1.2 [41, 25] RGB-D 87.2 90.1 88.7 9 60.8 64.2 62.5 9
-hccepose RGB 81.4 94.8 88.1 10 53.1 64.5 58.8 18 ( $\downarrow$8)
-hipose-cvpr24 [24] RGB-D 86.6 88.8 87.7 11 58.0 60.8 59.4 15 ( $\downarrow$4)
-cosypose [21] RGB 83.6 90.7 87.2 12 57.2 64.0 60.6 13 ( $\downarrow$1)
-pfa RGB-D 85.5 87.7 86.6 13 59.2 62.0 60.6 12 ( $\uparrow$1)
-hccepose RGB 79.2 93.2 86.2 14 51.4 63.1 57.2 20 ( $\downarrow$6)
-zebraposesa [37] RGB 79.1 93.0 86.1 15 51.8 63.4 57.6 19 ( $\downarrow$4)
-gdrnppde v1.1 [41] RGB-D 78.4 92.2 85.3 16 69.9 82.9 76.4 7 ( $\uparrow$9)
-leroy-fuseocclu-rgb RGB-D 78.0 91.7 84.8 17 70.4 83.3 76.9 6 ( $\uparrow$11)
-gdpnpp v1.0 [41, 25] RGB-D 83.4 85.6 84.5 18 57.5 60.2 58.9 17 ( $\uparrow$1)
-surfemb [6] RGB-D 82.9 85.9 84.4 19 75.8 78.7 77.2 5 ( $\uparrow$14)
-poseio RGB-D 81.7 86.0 83.8 20 69.3 73.6 71.4 8 ( $\uparrow$12)
+BOP [42] (object-wise annotations) BOP-Distrib (our image-wise annotations)
+BOP Contenders $\mathbf{MSSD}$ $\mathbf{MSPD}$ $\mathbf{Mean}$ Rank $\mathbf{MSSD}$ $\mathbf{MSPD}$ $\mathbf{Mean}$ $\mathbf{Loss}$ Rank
+T-LESS [13] BOP Challenge 2023 [17]: pose estimation of unseen objects
+foundpose [37] 55.0 62.3 58.6 1 45.5 52.1 48.8 -9.8 3( $\downarrow$-2)
+gigaposemegapose-5-hypothesis [35, 24] 54.3 62.0 58.1 2 45.3 52.3 48.8 -9.3 2
+gigaposemegapose-1-hypothesis [35, 24] 51.5 59.2 55.3 3 38.5 45.1 41.8 -13.5 9( $\downarrow$-6)
+genflow-multihypo16-rgb [32] 50.9 57.9 54.4 4 45.9 52.4 49.1 -5.3 1( $\uparrow$3)
+genflow-multihypo-rgb [32] 50.4 57.3 53.9 5 44.9 51.4 48.1 -5.8 4( $\uparrow$1)
+genflow-multihypo16 [32] 52.8 54.8 53.8 6 46.7 48.6 47.7 -6.1 5( $\uparrow$1)
+genflow-multihypo [32] 51.6 53.6 52.6 7 45.2 47.2 46.2 -6.4 6( $\uparrow$1)
+foundpose [37] 49.1 55.6 52.4 8 37.7 43.4 40.6 -11.8 11( $\downarrow$-3)
+foundpose [37] 49.1 55.6 52.3 9 37.7 43.4 40.5 -11.8 12( $\downarrow$-3)
+cnos-fastsammegapose-multihyp-10 [34, 24] 48.5 55.9 52.2 10 40.6 47.1 43.8 -8.4 7( $\uparrow$3)
+cnos-fastsammegapose-multihyp [34, 24] 48.4 55.6 52.0 11 40.5 47.1 43.8 -8.2 8( $\uparrow$3)
+RGB cnos-fastsammegapose-multihyp-teaserpp [34, 24] 48.4 51.3 49.9 12 40.3 43.0 41.7 -8.2 10( $\uparrow$2)
+foundationpose [46] 62.9 64.1 63.5 1 56.7 58.0 57.4 -6.1 1
+gigaposegenflowkabsch [35] 56.3 58.6 57.5 2 49.4 51.7 50.6 -6.9 2
+sam6d [26] 52.9 54.5 53.7 3 39.7 41.2 40.4 -13.3 3
+sam6d [26] 52.1 53.5 52.8 4 38.3 39.7 39.0 -13.8 5( $\downarrow$-1)
+sam6d-fastsam [26] 49.9 51.1 50.5 5 37.9 39.1 38.5 -12.0 6( $\downarrow$-1)
+sam6d-cnosmask [26] 50.0 50.9 50.4 6 39.3 40.3 39.8 -10.6 4( $\uparrow$2)
+sam6d-cnosfastsam [26] 49.3 50.5 49.9 7 37.2 38.4 37.8 -12.1 7
+sam6d-zeropose [26] 44.4 45.6 45.0 8 33.5 34.8 34.2 -10.8 8
+RGB-D zeropose-multi-hypo-refinement-defaultseg [4] 41.5 42.0 41.8 9 29.4 30.1 29.7 -12.1 9
+T-LESS [13] BOP Challenge 2022 [42]: pose estimation
+hccepose-2024-pbr 82.9 95.8 89.4 1 53.8 65.2 59.5 -29.9 5( $\downarrow$-4)
+hcceposeefficientnet-b4-default-2d-bbox 81.4 94.8 88.1 2 53.1 64.5 58.8 -29.3 6( $\downarrow$-4)
+cosypose-eccv20-syntreal-8views [23] 83.6 90.7 87.2 3 57.2 64.0 60.6 -26.6 4( $\downarrow$-1)
+hccepose-default-2d-bbox 79.2 93.2 86.2 4 51.4 63.1 57.2 -29.0 8( $\downarrow$-4)
+gdrnppdet-pbrrealgenflow-multihypo-rgb [44, 28] 78.4 92.2 85.3 5 69.9 82.9 76.4 -8.9 2( $\uparrow$3)
+leroy-fuseocclu-rgb 78.0 91.7 84.8 6 70.4 83.3 76.9 -7.9 1( $\uparrow$5)
+gpose2023-rgb [48] 76.6 92.9 84.7 7 50.2 64.0 57.1 -27.6 9( $\downarrow$-2)
+gdrnpp-pbr-rgb-mmodel [44, 28] 76.3 92.4 84.4 8 50.0 63.6 56.8 -27.6 10( $\downarrow$-2)
+gdrnpp-pbrreal-rgb-mmodel [44, 28] 76.0 91.3 83.6 9 48.8 61.7 55.2 -28.4 13( $\downarrow$-4)
+cosypose-eccv20-syntreal-4views [23] 79.5 86.4 83.0 10 52.7 59.3 56.0 -27.0 12( $\downarrow$-2)
+mrc-net [25] 78.5 87.1 82.8 11 52.3 60.3 56.3 -26.5 11
+gdrnppdet-pbrrealmegapose-multihyp [44, 28] 74.0 87.4 80.7 12 61.6 73.8 67.7 -13.0 3( $\uparrow$9)
+RGB sc6d [3] 72.1 85.3 78.7 13 52.6 64.1 58.3 -20.4 7( $\uparrow$6)
+gpose2023 [48] 92.1 94.6 93.4 1 86.5 89.1 87.8 -5.6 1
+gdrnppv2-rgbd-pbrreal [44, 28] 90.2 92.6 91.4 2 84.8 87.2 86.0 -5.4 3( $\downarrow$-1)
+gpose2023-pbr [48] 90.3 92.6 91.4 3 84.7 87.2 86.0 -5.4 2( $\uparrow$1)
+modalocclusion-rgbd 89.7 92.9 91.3 4 83.2 86.4 84.8 -6.5 4
+hcceposebf-2024-ref 89.0 91.7 90.3 5 71.7 74.7 73.2 -17.1 6( $\downarrow$-1)
+gdrbpp-pbrreal-rgbd-mmodel-fast v1.4 [44, 28] 88.7 91.5 90.1 6 58.1 61.9 60.0 -30.1 13( $\downarrow$-7)
+gdrnpp-pbrreal-rgbd-mmodel v1.3 [44, 28] 88.4 90.9 89.6 7 60.6 63.7 62.2 -27.4 10( $\downarrow$-3)
+gdrnpp-pbrreal-rgbd-smodel v1.2 [44, 28] 87.1 90.3 88.7 8 57.5 61.2 59.3 -29.4 15( $\downarrow$-7)
+zebraposesat-effnetb4-refineddefdet-2023 [40] 87.3 90.2 88.7 9 60.2 63.4 61.8 -26.9 11( $\downarrow$-2)
+gdrnpp-pbr-rgbd-mmodel v1.1 [44, 28] 87.2 90.1 88.7 10 60.8 64.2 62.5 -26.2 9( $\uparrow$1)
+hipose-cvpr24 [27] 86.6 88.8 87.7 11 58.0 60.8 59.4 -28.3 14( $\downarrow$-3)
+defaultdetection-pfa-mixpbr-rgb-d 85.5 87.7 86.6 12 59.2 62.0 60.6 -26.0 12
+zebraposesat-effnetb4defdet-2023 [40] 79.1 93.0 86.1 13 51.8 63.4 57.6 -28.5 17( $\downarrow$-4)
+zebraposesat-effnetb4pbr-only-defdet-2023 77.7 92.0 84.9 14 50.7 62.8 56.8 -28.1 18( $\downarrow$-4)
+gdpnpp-pbrreal-rgbd-mmodel-officialdet v1.0 [44, 28] 83.4 85.6 84.5 15 57.5 60.2 58.9 -25.6 16( $\downarrow$-1)
+surfemb-pbr-rgbd-lin [8] 82.9 85.9 84.4 16 75.8 78.7 77.2 -7.2 5( $\uparrow$11)
+RGB-D poseio 81.7 86.0 83.8 17 69.3 73.6 71.4 -12.4 8( $\uparrow$9)
+YCB-V [47] BOP Challenge 2022 [42]: pose estimation
+cosypose-eccv20-syntreal-8views [23] 88.5 88.0 88.2 1 83.3 82.4 82.9 -5.3 2( $\downarrow$-1)
+cosypose-eccv20-syntreal-4views [23] 86.9 86.6 86.8 2 82.1 81.3 81.7 -5.1 4( $\downarrow$-2)
+mrpe-pbrreal-rgb-smodel 85.7 87.3 86.5 3 83.1 84.4 83.7 -2.8 1( $\uparrow$2)
+RGB magic 85.9 85.4 85.6 4 82.9 82.3 82.6 -3.0 3( $\uparrow$1)
+gdrnppv2-rgbd-pbrreal [44, 28] 96.3 92.7 94.5 1 94.2 90.2 92.2 -2.3 1
+gpose2023 [48] 96.2 92.6 94.4 2 94.2 90.2 92.2 -2.2 2
+gdrnpp-pbrreal-rgbd-mmodel [44, 28] 95.7 91.9 93.8 3 90.1 86.2 88.1 -5.7 3
+RGB-D dsgc6d 93.0 90.1 91.5 4 87.5 85.2 86.4 -5.1 4
 T-LESS [13] Best mode of pose distribution methods
-SpyroPose [7] RGB 61.2 75.3 68.3 1 57.8 71.6 64.7 1
-LiePose Diffusion [17] RGB 60.1 92.2 76.1 2 48.3 76.1 62.2 2
-T-LESS [11] Challenge: pose estimation of unseen objects
-foundationpose [43] RGB-D 62.9 64.1 63.5 1 56.7 58.0 57.4 1
-foundpose [34] RGB 55.0 62.3 58.6 2 45.5 52.1 48.8 5 ( $\downarrow$3)
-gigaposemegapose [32, 22] RGB 54.3 62.0 58.1 3 45.3 52.3 48.8 4 ( $\downarrow$1)
-gigaposegenflowkabsch [32, 29] RGB-D 56.3 58.6 57.5 4 49.4 51.7 50.6 2 ( $\uparrow$2)
-gigaposemegapose [32] RGB 51.5 59.2 55.3 5 38.5 45.1 41.8 11 ( $\downarrow$6)
-genflow-multihypo [29] RGB 50.9 57.9 54.4 6 45.9 52.4 49.1 3 ( $\uparrow$3)
-genflow-multihypo [29] RGB 50.4 57.3 53.9 7 44.9 51.4 48.1 6 ( $\uparrow$1)
-genflow-multihypo16 [29] RGB-D 52.8 54.8 53.8 8 46.7 48.6 47.7 7 ( $\uparrow$1)
-sam6d [23] RGB-D 52.9 54.5 53.7 9 39.7 41.2 40.4 15 ( $\downarrow$6)
-sam6d [23] RGB-D 52.1 53.5 5208 10 38.3 39.7 39.0 18 ( $\downarrow$8)
-genflow-multihypo [29] RGB-D 51.6 53.6 52.6 11 45.2 47.2 46.2 8 ( $\uparrow$3)
-foundpose [43] RGB 49.1 55.6 52.4 12 37.7 43.4 40.6 13 ( $\downarrow$1)
-foundpose [43] RGB 49.1 55.6 52.3 13 37.7 43.4 40.5 14 ( $\downarrow$1)
-cnos-fastsam10 [31] RGB 48.5 55.9 52.2 14 40.6 47.1 43.8 10 ( $\uparrow$4)
-Table 1: Results of BOP 2023 on T-LESS [11].
-We report the top contenders results of the pose estimation challenge (2023 edition) evaluated with $\mathbf{MSSD}$ and $\mathbf{MSPD}$ on both the official ground truth and our ground truth.
+RGB LiePose Diffusion [19] 60.1 92.2 76.1 1 48.3 76.1 62.2 -13.9 2( $\downarrow$-1)
+SpyroPose [9] 61.2 75.3 68.3 2 57.8 71.6 64.7 -3.6 1( $\uparrow$1)
+Table 1: Results of BOP 2023 on T-LESS [13] and YCB-V [47] with our annotations.
+We report the top contenders results of the pose estimation of unseen objects challenge and the pose estimation challenge evaluated with $\mathbf{MSSD}$ and $\mathbf{MSPD}$ on both the official ground truth (object-wise) and our ground truth (image-wise).
 We rank the methods with the mean of $\mathbf{MSSD}$ and $\mathbf{MSPD}$ (similar to BOP, except we exclude $\mathbf{VSD}$ as BOP starts to abandon it) and re-rank them with the mean of $\mathbf{MSSD}$ and $\mathbf{MSPD}$ computed on our ground truth distributions.
-This produces drastic changes in the rankings.
+The drop in the scores highlighted by the $\mathbf{Loss}$ column ($\mathbf{Loss}=\mathbf{Mean}_{\mathbf{BOP-Distrib}}-\mathbf{Mean}_{\mathbf{BOP}}$) produces drastic changes in the rankings.
 We also show the same scores for the best mode of distribution estimation methods, however, both of them use ground truth bounding boxes instead of a detector, so direct comparison would be unfair.
-Finally, we apply the same study to the pose estimation of unseen objects challenge.
-Similarly, the rankings change.
 This highlights the importance to consider image level symmetries as we proposed, to evaluate accurately pose estimation methods.
-(names of the methods are automatically processed from BOP csv result files)
+We also processed YCB-V.
+The $\mathbf{Loss}$ is much lower as there are fewer symmetrical objects.
+Names of the methods are automatically processed from BOP csv result files.
 5.3 Single Pose Estimation Evaluation
@@
 We considered top contenders of the BOP challenge for the pose estimation and pose estimation for unseen objects tasks.
-We also included two methods for the multi-modal pose distribution estimation task [7, 17] for which we kept only the mode with highest confidence mode.
+We also included two methods for the multi-modal pose distribution estimation task [9, 19] for which we kept only the mode with highest confidence.
 We consider their full predicted distributions in the next subsection.
-Since these methods are not part of BOP challenge, we retrained them on the T-LESS train set including both photo-realistic synthetic images and real images.
+Since these methods are not part of the BOP challenge, we retrained them on the T-LESS train set.
 Experimental protocol.
 We evaluate the current (July 2024) top contenders of the BOP Challenge on the T-LESS dataset [13] on the pose estimation task and the pose estimation of unseen object task together with [9, 19].
+We conduct a similar evaluation with YCB-V [47].
 This evaluation is based directly on the raw pose results provided by the contenders and the official evaluation scripts of BOP challenges, both being publicly available on the Challenge website.
@@
 Our method ranking is based on the mean of Recalls computed on MSPD and MSSD accuracies.
-As discussed in Section 2, we do not consider VSD metric as it is now omitted for new tasks in BOP.
 Results.
-Table 1 reports the results of our evaluations.
-We can observe a large impact of the ground truth on the methods ranking since only the first place in the ranking remains unchanged.
-For example, regarding the ranking of the pose estimation task, the method Surfemb [6] moves from the 19th place with the original ground truth annotation to the 5th place with our annotation, gaining 14 places.
-For pose estimation of unseen objects, the difference of ranking can reach 8 places.
+Table 1 reports the results.
+We can observe a large impact of the ground truth on the methods results (see Loss column) and ranking.
+In terms of ranking, the well established pose estimation task face the biggest changes, with gdrnppdet-pbrrealmegapose-multihyp moving from the 12th to the 3rd position for RGB methods and surfemb-pbr-rgbd-lin [8] moving from the 16th to the 5th place for the RGB-D methods.
+For pose estimation of unseen objects, the difference of ranking can reach 6 places.
 We observe large impacts on the recall performances as well.
-Whereas all the 20 first contenders of pose estimation task have a mean recall up to 80% and even up to 90% for 5 methods with the original annotations, only the 4 methods exceed this score with our annotation, with a majority of method having less than 65%.
-Similarly, all the 14 top contenders for the pose estimation of unseen objects task have a mean recall that exceed 50% with original annotation whereas only the best one exceed this score once evaluated with our annotations.
-It is also interesting to note that the ranking order of variants 1.3, 1.4 and 1.5 of GDRNPP (based on different object detectors, training augmentations, and refinement methods) is changed when evaluated with our ground truth.
+Whereas 29 of the 30 contenders of pose estimation task have a mean recall up to 80% and even up to 90% for 6 methods with the original annotations, only 4 methods exceed this score with our annotations, with a majority of methods having less than 65%.
+Similarly, most of the contenders for the pose estimation of unseen objects task have a mean recall that exceeds 50% with the original annotations whereas only the 2 best RGB-D methods exceed this score once evaluated with our annotations.
+As shown in Figure 1 (Case 3), the image-wise annotation rejects poses when the image is not ambiguous.
+The Loss is explained by inaccurate pose estimates that were mistakenly validated by previous object-wise annotations (see supplementary material for visualizations).
+It is also interesting to note that the ranking order of variants v1.2, v1.3 and v1.4 of GDRNPP RGB-D (based on different object detectors, training augmentations, and refinement methods) is changed when evaluated with our ground truth.
 It underlines that inaccurate ground truth annotation can induce erroneous conclusions in ablation studies.
 5.4 Pose Distribution Evaluation
-In this experiment, we provide the first evaluation of 6D pose distribution estimation methods on a dataset of real images.
+We provide now the first evaluation of 6D pose distribution estimation methods on a dataset of real images.
 Experimental protocol
-Similarly to Section 5.3, the methods were trained on the train set of T-LESS that includes both photo-realistic synthetic images and real images.
-The ground truth used for this dataset is the same than the one of Section 5.3 and is constituted of a discreet set of 6D pose obtained with the 6D pose annotation process introduced in Section 3.
+Similarly to Section 5.3, the methods were trained on the train set of T-LESS.
+The ground truth used for this dataset is the same than the one of Section 5.3 and is constituted of a discrete set of 6D pose obtained with the 6D pose annotation process introduced in Section 3.
 The performances of the methods are evaluated in Precision and Recall, using accuracy measures MPD and MSD as introduced in Section 4.
@@
 Results
-We evaluated SpyroPose [7] and LiePose-Diffusion [17] since they provide 6D pose distribution, unlike [30, 16] that provide distribution only over the rotation.
-Quantative results are reported on Table 2 whereas qualitative results are provided in the supplementary material.
-It appears that LiePose-Diffusion outperforms SpyroPose. Without surprise, performances decrease when MSD accuracy is used since those two methods use RGB images only.
-Methods $\mathbf{P_{MPD}}$ $\mathbf{R_{MPD}}$ $\mathbf{P_{MSD}}$ $\mathbf{R_{MSD}}$
-SpyroPose [7] 69.9 69.9 55.2 65.9
-LiePose-Diff [17] 75.4 94.1 48.1 85.3
-Table 2: Comparisons of pose distribution estimation methods on T-LESS using our ground truth pose distributions.
-So far, these methods were evaluated quantitatively on synthetic images only and only qualitatively on real images.
+We evaluated SpyroPose [9] and LiePose-Diffusion [19] (either PBR + real images and PBR only to compare with SpyroPose) since they provide 6D pose distributions, unlike [33, 18] that provide distributions only over the rotation.
+To show genericity, we also evaluated the top 3 RGB methods of single pose estimation task (with image-wise annotations).
+Their distribution is considered as a Dirac distribution on the single estimate.
+Results are given on the complete dataset, and the subset of objects with cylindrical symmetry (1-4, 13-18, 24 and 30, as illustrated in supplementary material).
+Quantitative results are reported on Table 2 whereas qualitative results are provided in the supplementary material.
+It appears that LiePose-Diffusion outperforms SpyroPose, except for the Precision MSD.
+Lower accuracy are obtained with MSD error since those methods use RGB images only and no depth.
+Single pose methods outperform distribution methods in terms of Precision, as they are optimized for that.
+Yet, distribution methods produce much stronger Recalls, even more on only symmetrical objects, highlighting their ability to retrieve multiple meaningfull poses from the targeted image-wise distribution.
+Methods Train $\mathbf{P_{MSD}}$ $\mathbf{R_{MSD}}$ $\mathbf{P_{MPD}}$ $\mathbf{R_{MPD}}$
+T-LESS [13] Complete dataset
+Distrib SpyroPose [9] PBR 32.8 48.1 55.9 55.5
+LiePose-Diffusion [19] PBR 24.6 71.4 61.2 89.5
+LiePose-Diffusion [19] PBR + real 29.9 75.4 68.4 90.6
+leroy-fuseocclu-rgb PBR + real 55.4 35.7 80.9 54.2
+gdrnppdet-pbrrealgenflow-multihypo-rgb [44, 28] PBR + real 52.4 34.9 77.7 52.2
+Single cosypose-eccv20-syntreal-8views [23] PBR + real 43.5 25.7 53.1 30.9
+T-LESS [13] Only symmetrical objects
+Distrib SpyroPose [9] PBR 31.5 42.7 70.5 60.8
+LiePose-Diffusion [19] PBR 18.3 61.6 63.6 87.3
+LiePose-Diffusion [19] PBR + real 24.1 67.0 70.3 88.1
+leroy-fuseocclu-rgb PBR + real 52.4 25.7 86.1 45.3
+gdrnppdet-pbrrealgenflow-multihypo-rgb [44, 28] PBR + real 45.8 20.8 76.5 37.8
+Single cosypose-eccv20-syntreal-8views [23] PBR + real 29.1 6.5 40.0 10.5
+Table 2: Comparison of pose distribution estimation methods on T-LESS using our ground truth pose distributions.
 6 Limitations
@@
 We selected their values carefully.
-We let the reader check the visualizations in the supplementary material.
+We invite the reader to check the visualizations in the supplementary material.
 Our annotation method relies on a geometric analysis of the scene.
 It is possible that the sensor resolution, sensor noise, field of view, or motion blur may also affect the visibility of the disambiguating parts.
-Considering these effects would improve even more the ground truth.
+Considering these effects would improve the ground truth even further.
 7 Conclusion
-For simplicity, most of the 6D pose estimation benchmarks rely on a single 6D pose annotation per image, the latter being completed with a per-object symmetry pattern to transform this unique pose into a distribution.
+For simplicity, most of the 6D pose estimation benchmarks rely on a single 6D pose annotation per image, completed by a per-object symmetry pattern to transform this unique pose into a distribution.
 We argued that ignoring the per-image nature of the symmetry pattern is prone to bias in the resulting distribution ground truth and in the resulting performances evaluations.
@@
 Moreover, when using this ground truth to re-evaluate current state-of-the-art methods, we showed that the ranking of these methods changes drastically.
-We also introduced metrics to evaluate methods that estimate a pose distribution and provided the first evaluation of such methods on real data.
+We also introduced metrics to evaluate methods that estimate a pose distribution and provided their first evaluation on real data.
 We believe that these results should encourage the 3D vision community to adopt a new outlook on these methods and how well they work in practice, and should guide future research.
@@
 Only Object 18 is perfectly symmetrical and our method retrieves the same poses as the original annotations.
-For the other objects, in particular the objects with complex symmetry patterns like the fist 4 objects.
+For the other objects, in particular the objects with complex symmetry patterns like the first 4 objects.
 Our annotations significantly change the original annotations.

### v3

@@
 Abstract
-6D pose estimation aims at determining the pose of the object that best explains the camera observation.
-The unique solution for a non-symmetrical object can turn into a multi-modal pose distribution for a symmetrical object or when occlusions of symmetry-breaking elements happen, depending on the viewpoint.
+6D pose estimation aims at determining the object pose that best explains the camera observation.
+The unique solution for non-ambiguous objects can turn into a multi-modal pose distribution for symmetrical objects or when occlusions of symmetry-breaking elements happen, depending on the viewpoint.
 Currently, 6D pose estimation methods are benchmarked on datasets that consider, for their ground truth annotations, visual ambiguities as only related to global object symmetries, whereas they should be defined per-image to account for the camera viewpoint.
-We thus first propose an automatic method to re-annotate those datasets with a 6D pose distribution specific to each image, taking into account the visibility of the object surface in the image to correctly determine the visual ambiguities.
+We thus first propose an automatic method to re-annotate those datasets with a 6D pose distribution specific to each image, taking into account the object surface visibility in the image to correctly determine the visual ambiguities.
 Second, given this improved ground truth, we re-evaluate the state-of-the-art single pose methods and show that this greatly modifies the ranking of these methods.
 Third, as some recent works focus on estimating the complete set of solutions, we derive a precision/recall formulation to evaluate them against our image-wise distribution ground truth, making it the first benchmark for pose distribution methods on real images.
-We will make our annotations for the T-LESS dataset and our code publicly available.
 Figure 1: We provide for the first time 6D pose annotations in the form of a per-image object pose distribution.
@@
 Our method exploits the current annotations as a starting point, the ground truth visibility masks of the objects, and their 3D models.
-While we demonstrate it on T-LESS [13], our method is fully automatic and can be applied to other datasets as demonstrated on YCB-V [47].
+Its genericity is demonstrated on T-LESS [13] and YCB-V [48].
 We also re-evaluate classical metrics with these annotations to assess the performance of pose estimation methods that return a single pose estimate per image.
@@
   2. 2.
-    A comprehensive re-evaluation of 6-DOF single pose estimation methods related to the re-annotation of the T-LESS [13] ground-truth;
+    A comprehensive re-evaluation of 6-DOF single pose estimation methods related to the re-annotation of the T-LESS [13] and YCB-V [48] ground truths;
   3. 3.
@@
 Object 6D pose estimation aims at determining an object pose that best explains the camera observation.
-While the solution is unique when the object is non-symmetrical and not occluded, multiple equivalent possible solutions arise when the object is symmetrical or when the occlusions prevent the observation of symmetry-breaking elements.
+Multiple equivalent solutions arise when the object is symmetrical or when occlusions (external or self) prevent the observation of symmetry-breaking elements (e.g.
+for a cup [12, 30]).
 While the state of the art mostly focused on estimating a single pose [6, 39, 15, 24, 42, 45, 8, 41], some recent works focus on estimating the complete set of solutions [30, 37, 34, 23, 31, 18, 9, 19].
-As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are required, even for methods that return a single pose per image.
+As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are then required, even for methods that return a single pose per image.
 Datasets and 6D pose annotation.
 Various techniques have been introduced to annotate image datasets with their 6D poses.
-For synthetic datasets [2, 43], the ground truth pose is directly available, while, for real datasets, this pose is usually determined with the help of user interaction [7, 14, 5], markers [10, 11, 1], or a robotic arm [45, 31], and refined with ICP in the case of an RGB-D camera [7, 14, 45, 31].
+For synthetic datasets [2, 44], the ground truth pose is directly available, while, for real datasets, this pose is usually determined with the help of user interaction [7, 14, 5], markers [10, 11, 1], or a robotic arm [46, 32, 21], and refined with ICP in the case of an RGB-D camera [7, 14, 46, 32].
 However, such annotations only provide a single pose solution while multiple may exist.
@@
 This set of transforms, also called symmetries pattern, can then be pre-computed offline and applied to the initial solution to recover the whole set of solutions [16].
-For non-symmetrical objects, recovering the complete set of solutions is much more challenging since symmetries may arise from the non-visibility of disambiguating parts of the object, those non-visibilities being induced by the viewpoint (self-occlusion), by other elements of the scene (external occlusion).
+For non-symmetrical objects, recovering the complete set of solutions is much more challenging since symmetries may arise from the non-visibility of disambiguating parts of the object, those non-visibilities being induced by the viewpoint (self-occlusion), by other elements of the scene (external occlusion), or by the lack of image resolution (resolution occlusion).
 The transformation set that relates the different solutions is then specific to each image and cannot be pre-computed from the 3D object model.
 To our knowledge, no method has been proposed to determine this ground-truth transformation set per image.
-Instead, the current gold standard in 6D pose estimation benchmarks still consists in approximating this per-image transformation set with a unique global transformation set.
-The latter is usually computed with the same method than fully symmetric object with large enough tolerance to ignore small disambiguating elements (e.g.
-surface deviation tolerance of 1.5cm or 10% of the object diameter [16]).
+Instead, the current gold standard in 6D pose estimation benchmarks still consists in approximating this per-image transformation set with a unique global transformation set, usually computed with the same method than fully symmetric object with large enough tolerance to ignore small disambiguating elements (e.g.
+surface deviation tolerance of 1.5cm or 10% of the object diameter [16]), and cannot be used to evaluate properly multi-modal pose distribution estimates like [9, 19].
 Evaluation metrics.
@@
 robotics or Augmented Reality), deviations can be measured in 3D space (3D distance) or in the image space (reprojection error), and the registration error can be considered as the mean or maximal error over the object surface.
-In case of multi-valued ground-truths due to a symmetrical object or occlusion-induced symmetries, the accuracy corresponds to the minimal error with respect to the set of solutions [12, 47, 43].
+In case of multi-valued ground-truths due to a symmetrically tagged object, the accuracy corresponds to the minimal error with respect to the set of solutions [12, 48, 44].
 In practice, the most commonly used accuracy measures are :
   1. 1.
-    Average Distance metric (ADD), and variations for symmetrical objects (ADD-S [47], ADD-H [43]), target to measure mean Euclidean error between surface points of the object transformed with the estimated pose and the ground truth pose, but are replaced by MSSD now.
+    Average Distance metric (ADD) [11], and variations for symmetrical objects (ADD-S [48], ADD-H [44]), measure mean Euclidean error between estimated and ground truth surface points, but are replaced by MSSD.
   2. 2.
@@
   3. 3.
-    Maximum Symmetry-Aware Surface Distance (MSSD) [16] which is similar to ADD but considers the maximal error instead of the mean error and symmetry management.
-MSSD provides a 3D error, useful to robotics applications.
+    Maximum Symmetry-Aware Surface Distance (MSSD) [16] , similar to ADD, considers the maximal Euclidean error with symmetry management, and provides a 3D error, useful to robotics applications.
   4. 4.
-    Maximum Symmetry-Aware Projection Distance (MSPD) [16] measures the maximum reprojection error between the projection of the visible points with the ground-truth pose and their corresponding projection with the pose to evaluate and symmetry management.
-MSPD provides a reprojection error, useful to augmented reality applications.
+    Maximum Symmetry-Aware Projection Distance (MSPD) [16] measures the maximum reprojection error between projections of the ground truth and estimated surface points, with symmetry management, and provides an error useful for augmented reality.
 Based on these accuracy measures, and inspired from the evaluation of the detection methods, single 6D pose estimation methods are usually evaluated through Precision/Recall.
@@
 3 Method
-As shown in Figure 2, the probability of a pose explaining a given image of the object is related to the visibility of the different points on the surface of the object.
-More formally, given an image $I$ and a ground truth pose $P_{\text{GT}}$ for a target object in this image, the objective of our method is to recover the set ${\cal T}=\{T_{i}\in{\text{SE}}(3)\}$ of rigid transformations that transform the ground truth pose $P_{\text{GT}}$ into another pose that results in the same appearance of the object.
-In the following, we will call set ${\cal T}$ the symmetry pattern of image $I$.
-Depending on the geometry $M$ of the object and the set ${\cal V}(M)$ of its visible points, set ${\cal T}$ can be limited to the identity transform in the case of a non-ambiguous pose or can include multiple transformations in addition to the identity.
-As introduced in [12], ${\cal T}$ can be defined as partial $\epsilon$-symmetries of ${\cal V}(M)$:
+Consider a 3D object model $M$ observed from a pose $P_{\text{GT}}$.
+We call symmetry-pattern the set ${\cal T}=\{T_{i}\in{\text{SE}}(3)\}$ of rigid transformations that, once combined with $P_{\text{GT}}$, generate representations that are indistinguishable from the one generated by $P_{\text{GT}}$.
+Depending on the object geometry $M$ and the representation type, the set ${\cal T}$ can be limited to the identity transform or can include multiple transformations in addition to the identity.
+Previous work well studied the representation corresponding to a 3D surface, or a subset ${\cal V}(M)\in M$ of the 3D object surface.
+In such case, the symmetries pattern ${\cal T}({\cal V}(M))$ can be defined through $\epsilon$-symmetries  [12]:
 $\epsilon\text{-sym}({\cal V}(M),M)=\{T_{i}:d(T_{i}*{\cal V}(M),M)<\epsilon\}\>,$ (1)
-where $d$ measures the mean deviation between the points of ${\cal V}(M)$ and $M$, $\epsilon$ is the maximal surface deviation tolerance and * represents the application of the transformation $T_{i}$ onto the point cloud ${\cal V}(M)$.
-If the set ${\cal V}(M)$ of visible points is available, the symmetry pattern could in principle be approximated by sampling the rigid transformation space SE(3) and assessing whether the sample belongs to the $\epsilon$-symmetries.
-However, since sampling a 6D space becomes quickly computationally expensive, such annotation process was never used to our knowledge.
-In the following sections, we introduce solutions to reduce the computational cost (Section 3.1), how to introduce some tolerance in the evaluation of the $\epsilon$-symmetries (Section 3.2), and how to extend the annotation to textured objects (Section 3.3).
+where $\epsilon$ is the maximal deviation tolerance, * represents the application of the transformation $T_{i}$ onto the point set ${\cal V}(M)$ and $d$ measures the max deviation between the sets:
+$d(X,M)=\max_{x\in X}||x-\operatorname*{arg\,min}_{m\in M}(||x-m||_{2})||_{2}$.
+In this paper, we propose to consider the case of representation corresponding to an image $I$.
+The symmetries pattern ${\cal T}(I)$ associated to the image $I$ corresponds to a subset of the symmetries pattern ${\cal T}({\cal V}(M))$ associated to the visible part of the object ${\cal V}(M)$ in the image.
+Indeed, since $\epsilon$-symmetries work in 3D space, they do not take into account the potential self or external occlusions that can be induced by the projection process that produces the image.
+Consequently, to estimate ${\cal T}(I)$, our solution compute first ${\cal T}({\cal V}(M))$ then prune the pose visually-inconsistent (section 3.3).
+As illustrated in Figure 2, to achieve this estimation efficiently, our solution decompose the estimation of ${\cal T}({\cal V}(M))$ in a two step process: a first step of pre-computation (section 3.1) independent of ${\cal V}(M)$ and executed only once, and a second step that compute ${\cal T}({\cal V}(M))$ (section 3.2).
 3.1 Speeding up the computation of $\epsilon$-sym
-Computing $\epsilon$-symmetries for each instance of image of a dataset can quickly become time consuming since it implies to sample the 6D pose space around the ground truth pose.
-To accelerate the computation of the $\epsilon$-symmetries, we reformulate them to introduce image-independent terms that can be pre-computed.
+Computing ${\cal T}({\cal V}(M))$ for each instance in each image of a dataset can quickly become time consuming since it implies to sample the 6D pose space around the ground truth pose.
+This can explain why such annotation process was never used to our knowledge.
+We accelerate the $\epsilon$-symmetries computation by introducing terms independent from ${\cal V}(M)$ that can be pre-computed.
 First, by definition of the $\epsilon$-symmetries, for any subsets $V_{1}\subset{\cal V}(M)$ and $V_{2}\subset{\cal V}(M)$, we have:
@@
 $\epsilon\text{-sym}({\cal V}(M),M)=\bigcap_{v_{j}\in{\cal V}(M)}\epsilon\text{-sym}(v_{j},M)\>,$ (3)
-where $v_{j}$ is a 3D point and $\epsilon\text{-sym}(v_{j},M)$ is the symmetry pattern when the point $v_{j}$ is the only visible point, which we name elementary symmetry pattern of $v_{j}$.
-With such expression of the $\epsilon$-symmetries, it appears that elementary symmetry patterns can be pre-computed for the whole 3D object model $M$ since they are independent of the image.
+where $v_{j}$ is a 3D point and $\epsilon\text{-sym}(v_{j},M)$ is the symmetries pattern when the point $v_{j}$ is the only visible point, which we name elementary symmetries pattern of $v_{j}$.
+With such expression of the $\epsilon$-symmetries, it appears that elementary symmetries patterns can be pre-computed for the whole 3D object model $M$ since they are independent of the image.
 Only the set of visible points and the intersection of the corresponding pre-computed elementary symmetries patterns need to be computed per image.
@@
 $\forall i,\text{H}(T_{i})=\text{Card}\{v_{j}~{}|~{}v_{j}\in{\cal V}(M),T_{i}\in\epsilon\text{-sym}(v_{j},M)\}\>.$ (4)
-A strict intersection would only keep the $T_{i}$ with the maximum value of $\text{H}(T_{i})$ which is represented by $\text{H}(Id)$ as the identity of SE(3) and is always in the symmetry pattern.
+A strict intersection would only keep the $T_{i}$ with the maximum value of $\text{H}(T_{i})$ which is represented by $\text{H}(Id)$ as the identity of SE(3) and is always in the symmetries pattern.
 Our soft intersection tolerates a threshold and keeps all the $T_{i}$ such that $\text{H}(T_{i})>\text{H}(Id)-\tau$, where $\tau$ represents the minimum size (expressed in number of 3D points) of the symmetry-breaking elements.
+3.3 Refining $\epsilon$-sym: visual occlusions consideration
+[…]
+$Prune\big{(}D_{GT},D_{T_{i}},V_{GT},V_{T_{i}},\delta\big{)}=\\ \sum_{p\in V_{GT}\cup V_{T_{i}}}\begin{cases}0&\text{if}\;p\in V_{GT}\cap V_{T_{i}}\,\wedge\,\big{|}D_{GT}(p)-D_{T_{i}}(p)\big{|}<\delta\\ 1&\text{otherwise,}\end{cases}$ (5)
+where $D_{GT}$ and $V_{GT}$ are distance image and visible mask of the ground truth, $D_{T_{i}}$ and $V_{T_{i}}$ are their $T_{i}$ counterparts, with visible masks computed from sensor depth.
 3.4 Generalization to textured 3D objects
@@
 $\epsilon\text{-sym}({\cal V}(M),M)=\{T_{i}:d(T_{i}*{\cal V}(M),M)<\epsilon,\\ d_{\text{color}}(T_{i}*{\cal V}(M),M)<\zeta)\}\>,$ (6)
-where $d_{\text{color}}$ is the distance in color-space and $\zeta$ is the color deviation tolerance.
+where $\zeta$ is the color deviation tolerance and with:
+$d_{\text{color}}(X,M)=\max_{x\in X}d_{\text{col}}(x,\operatorname*{arg\,min}_{m\in M}(||x-m||_{2}))$,
+with $d_{\text{col}}(x,m)$ the color distance between vertices $x$ and $m$.
 4 Evaluating 6D Pose Distribution Predictions
-Currently, single pose estimation methods and distribution pose estimation methods are evaluated on different datasets (BOP challenge for single poses, SYMSOL for 6D pose distributions) with different metrics (registration error for single pose estimation, pose error for 6D pose distribution).
+Currently, single pose estimation methods and distribution pose estimation methods are evaluated on different datasets (BOP challenge for single poses, SYMSOL for 6D pose distributions) with different metrics (registration error for the first, pose error for the others).
 With our annotation method that provides the full set of 6D pose solutions even on real data, it becomes possible to evaluate both tasks on the same dataset.
@@
 First, whereas pose error is commonly used in 6D pose distribution evaluation process [34, 20, 9], we propose to use the registration error.
-The latter takes the object 3D shape into account, making the accuracy more meaningful for most of the applications.
+The latter takes the object 3D shape into account, making the accuracy more meaningful for most applications.
 It is also the type of error used by the current gold standard benchmark for single pose [43].
@@
 We define the precision in the case of pose distribution prediction as:
-$\mathbf{P_{d}}(\text{Est},\text{GT})=\\ \sum\limits_{P_{\text{Est}}^{j}\in Est}p(P_{\text{Est}}^{j})\left(\min_{P_{\text{GT}}\in\text{GT}}d(P_{\text{Est}}^{j},P_{\text{GT}})<\tau_{\mathbf{d}}\right)\>,$ (6)
-where Est and GT are respectively the estimated and the ground truth distributions to compare, and $p(\text{Est}_{i})$ represents the probability associated to the $i$-th element of the estimated distribution  [^2] If the method outputs solutions without assigning a probability to each pose, such as [19], a uniform probability over the solutions is used., $\mathbf{d}$ is the chosen registration distance, and $\tau_{\mathbf{d}}$ the associated threshold.
+$\mathbf{P_{d}}(\text{Est},\text{GT},\tau_{d})=\\ \sum\limits_{P_{\text{Est}}^{i}\in\text{Est}}p(P_{\text{Est}}^{i})\left(\min_{P_{\text{GT}}\in\text{GT}}\mathbf{d}(P_{\text{Est}}^{i},P_{\text{GT}})<\tau_{\mathbf{d}}\right)\>,$ (7)
+where Est and GT are respectively the estimated and the ground truth distributions to compare, and $p(P_{\text{Est}}^{i})$ represents the probability associated to the $i$-th element of the estimated distribution  [^2] If the method outputs solutions without assigning a probability to each pose, such as [19], a uniform probability over the solutions is used., $\mathbf{d}$ is the chosen registration distance, and $\tau_{\mathbf{d}}$ the associated threshold.
 Recall for pose distribution.
 Similarly to the precision, we define the recall in the case of pose distribution prediction as:
-$\mathbf{R_{d}}(\text{Est},\text{GT},\tau_{d})=\\ \sum\limits_{P_{\text{GT}}^{i}\in\text{GT}}\max\left(p(\hat{P}_{\text{Est}}^{i}),\frac{1}{\text{Card}(\text{GT})}\right)[d(\hat{P}_{\text{Est}}^{i},P_{\text{GT}}^{i})<\tau_{\mathbf{d}}],\\ \text{with }\hat{P}_{\text{Est}}^{i}=\operatorname*{arg\,min}_{P_{\text{Est}}\in\text{Est}}d(P_{\text{Est}},P_{\text{GT}}^{i})\>.$ (7)
+$\mathbf{R_{d}}(\text{Est},\text{GT},\tau_{d})=\\ \sum\limits_{P_{\text{GT}}^{i}\in\text{GT}}\min\left(p(\hat{P}_{\text{Est}}^{i}),\frac{1}{\text{Card}(\text{GT})}\right)[\mathbf{d}(\hat{P}_{\text{Est}}^{i},P_{\text{GT}}^{i})<\tau_{\mathbf{d}}],\\ \text{with }\hat{P}_{\text{Est}}^{i}=\operatorname*{arg\,min}_{P_{\text{Est}}\in\text{Est}}\mathbf{d}(P_{\text{Est}},P_{\text{GT}}^{i})\>.$ (8)
 The probability of $\hat{P}_{\text{Est}}^{i}$ is clamped to $\frac{1}{\text{Card}(\text{GT})}$ since ground truth poses are considered as equiprobable.
@@
 Since our annotation process and metrics are related to objects with intrinsic or occlusion-induced symmetries, we perform our evaluations on the T-LESS dataset.
-Indeed, among the datasets of the BOP challenge [17], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground-truth annotations: ITODD [7] and Homebrewed [21] feature symmetrical objects, with occlusion-induced symmetries, but do not publicly share their ground-truths.
+Indeed, among the datasets of the BOP challenge [17], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground-truth annotations: ITODD [7] and HB [22] feature symmetrical objects, with occlusion-induced symmetries, but do not publicly share their ground-truths.
 HOPE [44] and IC-BIN [5] have symmetrical shapes, but texture information disambiguates them completely.
 The first experiment focus on a qualitative evaluation of our annotation method (Section 5.2).
-The second evaluates the impact of our re-annotation of T-LESS Dataset on the performance evaluation of the best performer methods of the BOP benchmark [17] on single pose estimation  (the Multiple instances of Multiple known objects (MiMo) task of BOP) (Section 5.3).
-The last experiment evaluates for the first time methods of 6D pose distribution estimation on a non-synthetic dataset with proper annotations and metrics (Section 5.4).
+The second evaluates the impact of our re-annotation of T-LESS and YCB-V datasets on the performance evaluation of the methods of the BOP benchmark [17] on single pose estimation  (Section 5.3).
+The last experiment evaluates for the first time 6D pose distribution methods on a non-synthetic dataset with proper annotations and metrics (Section 5.4).
 5.1 Implementation details
 […]
-The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object.
-For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 2.5$\times$2.5mm2.
+Surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object.
+For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating size of roughly 2.5$\times$2.5mm2, $\delta=5mm$ and $\tau_{pix}=30pix$.
 5.2 Qualitative validation of the method
@@
 5.3 Single Pose Estimation Evaluation
-Since May 2023, raw results of BOP Challenge [17] public submissions have become publicly available to allow in-depth analysis.
-In this experiment, we use them to compare the performance evaluations of state-of-the-art methods on T-LESS for both the original 6D annotations and our annotations.
+Since May 2023, raw results of BOP Challenge [17] submissions have become publicly available to allow in-depth analysis.
+This experiment reprocess these results to compare the performance evaluations of state-of-the-art methods for both the original 6D annotations and our annotations.
 We considered top contenders of the BOP challenge for the pose estimation and pose estimation for unseen objects tasks.
-We also included two methods for the multi-modal pose distribution estimation task [9, 19] for which we kept only the mode with highest confidence.
-We consider their full predicted distributions in the next subsection.
-Since these methods are not part of the BOP challenge, we retrained them on the T-LESS train set.
+We also included two methods for the multi-modal pose distribution estimation task [9, 19] for which we kept only the highest confidence mode.
+Since these methods are not part of the BOP challenge, we retrained them.
 Experimental protocol.
@@
 This evaluation is based directly on the raw pose results provided by the contenders and the official evaluation scripts of BOP challenges, both being publicly available on the Challenge website.
-Following the BOP Challenge, the evaluation is based exclusively on recall.
 Our method ranking is based on the mean of Recalls computed on MSPD and MSSD accuracies.
@@
 We can observe a large impact of the ground truth on the methods results (see Loss column) and ranking.
-In terms of ranking, the well established pose estimation task face the biggest changes, with gdrnppdet-pbrrealmegapose-multihyp moving from the 12th to the 3rd position for RGB methods and surfemb-pbr-rgbd-lin [8] moving from the 16th to the 5th place for the RGB-D methods.
+In terms of ranking, the well established pose estimation task faces the biggest changes, with gdrnppdet-pbrrealmegapose-multihyp moving from the 12th to the 3rd position for RGB methods and surfemb-pbr-rgbd-lin [8] moving from the 16th to the 5th place for the RGB-D methods.
 For pose estimation of unseen objects, the difference of ranking can reach 6 places.
@@
 As shown in Figure 1 (Case 3), the image-wise annotation rejects poses when the image is not ambiguous.
-The Loss is explained by inaccurate pose estimates that were mistakenly validated by previous object-wise annotations (see supplementary material for visualizations).
+The Loss is explained by inaccurate pose estimates that were mistakenly validated by previous object-wise annotations (see supplementary materials for visualizations).
 It is also interesting to note that the ranking order of variants v1.2, v1.3 and v1.4 of GDRNPP RGB-D (based on different object detectors, training augmentations, and refinement methods) is changed when evaluated with our ground truth.
@@
 We provide now the first evaluation of 6D pose distribution estimation methods on a dataset of real images.
-Experimental protocol
+Experimental protocol.
 Similarly to Section 5.3, the methods were trained on the train set of T-LESS.
@@
 The performances of the methods are evaluated in Precision and Recall, using accuracy measures MPD and MSD as introduced in Section 4.
-More details on methods processing and results formatting are given in the supplementary material.
-Results
+More details on methods and results processing are given as supplementary material.
+Results.
 We evaluated SpyroPose [9] and LiePose-Diffusion [19] (either PBR + real images and PBR only to compare with SpyroPose) since they provide 6D pose distributions, unlike [34, 18] that provide distributions only over the rotation.
@@
 6 Limitations
-While our solution improves the pose annotations, the resulting pose distribution depends on some parameters (tolerance factors, sampling of the SE(3) space, etc.).
-We selected their values carefully.
-We invite the reader to check the visualizations in the supplementary material.
 Our annotation method relies on a geometric analysis of the scene.
@@
 Supplementary Material
-1 Analysis of our BOP-D Annotations
-1.1 Additionnal BOP-D Ground Truths Visualizations
+1 Pseudo-code of Per-image Ground Truth Pose Annotation
+In this section, we give the key elements of our annotation method with the following pseudo-codes.
+1# Input: CAD model, candidate transformation set, threshold
+2# Output: per point epsilon sym set
+3def EpsSym(ModelCAD, TransformSet, $\epsilon$, $\zeta$, resolution):
+4 SampledModel = uniform_resampling(ModelCAD, resolution)
+5 for point in SampledModel:
+6 for T in TransformSet:
+7 # Texture-less case
+8 if testColor == False:
+9 if knnSearch(T*point, SampledModel, $\epsilon$) == True:
+10 EpsSym[point].append(T)
+11 # Textured case
+12 else:
+13 if knnSearch(T*point, SampledModel, $\epsilon$) == True:
+14 if $d_{Color}$(point, T*point) < $\zeta$:
+15 EpsSym[point].append(T)
+16
+17 return EpsSym, SampledModel
+[…]
+The geometric ($d_{Geom}$) and colorimetric ($d_{Color}$) distances are implemented as follows:
+$d_{Geom}(x,m)=||x-m||_{2}$ and
+$d_{Color}(x,m)<\zeta\Leftrightarrow\min\Big{(}|h(x)-h(m)|,\\ |h(x)-h(m)-360|,|h(x)-h(m)+360|\Big{)}<\zeta_{h},\\ \text{AND~{}}|s(x)-s(m)|<\zeta_{s},\\ \text{AND~{}}|v(x)-v(m)|<\zeta_{v}.$
+with $h(.),s(.),v(.)$ being hue, saturation and value of the point, and $\zeta_{h},\zeta_{s},\zeta_{v}$, the respective hue, saturation and value thresholds.
+1# Input: SampledModel, EpsSym, PoseGT, MaskVisib, $\tau$
+2# Output: EpsSymImage
+3def SoftIntersection(SampledModel, EpsSym, PoseGT, MaskVisib, $\tau$):
+4 # Count the vertices that vote for a transform
+5 for point in SampledModel:
+6 if K*[$R_{gt}$, $T_{gt}$]*point in mask_visib:
+7 for T in EpsSym[point]:
+8 H[T]++
+9 Sort(H)
+10 # Count the vertices that vote for a transform
+11 for i in size(H):
+12 if H[0]-H[i] < $\tau$:
+13 EpsSymImage.append(i)
+14
+15 return EpsSymImage
+Listing 2: Image annotation, pseudo-code of equation (4) from section 3.2.
+1# Input: ModelCAD, EpsSymImage, PoseGT, SensorDepth, $\delta$
+2# Output: EpsSymImageGlobal
+3def PostProcessAnnotateImage(ModelCAD, EpsSymImage, PoseGT, SensorDepth, $\delta$, threshold):
+4 depthGT = render(ModelCAD, PoseGT)
+5 distGT = depthToDistImage(depthGT)
+6 visibleMaskGT = generateMask(depthGT, SensorDepth)
+7 for Ti in EpsSymImage:
+8 depthEst = render(ModelCAD, Ti)
+9 distEst = depthToDistImage(depthEst)
+10 visibleMaskEst = generateMask(depthEst, SensorDepth)
+11 maskIntersection = intersection(visibleMaskGT, visibleMaskEst)
+12 maskUnion = union(visibleMaskGT, visibleMaskEst)
+13 nbPixelsOutlier += maskUnion-maskIntersection
+14 nbPixelsOutlier += abs(distGt - distEst)[maskIntersection] >= $\delta$
+15 if nbPixelsOutlier < threshold:
+16 EpsSymImageGlobal.append(Ti)
+17
+18 return EpsSymImageGlobal
+Listing 3: Image annotation depth post-processing, pseudo-code of equation (4) from section 3.3, based on VSD implementation from BOP toolkit.
+Section 3.3 does not use the VSD metric, as VSD is normalized by the number of pixels in the union of the visible masks.
+In our case, we need an absolute score and not a relative one, so that the counting of outliers has a metric meaning and can represent the size of a minimal disambiguating element.
+2 Analysis of our BOP-Distrib Annotations
+2.1 Accuracy of the annotations
+Figure 4 illustrate renderings of all poses of a symmetries pattern.
+Figure 4: Variation between poses in a single mode symmetries pattern.
+We display the renderings of the symmetries pattern of object 2 in image 7, scene 1.
+Although the SO(3) visualization may show spread points, these rotations result in similar poses.
+2.2 False visible points and false occluded points detected by the pruning stage
+[…]
+[…]
+[…]
+2.3 Additionnal BOP-Distrib Ground Truths Visualizations
 We first provide more visualizations for qualitative appreciation of the new ground truth annotations accuracy in Figure 11.
-These images are taken from a video compilation of all ground truth annotations sorted by object identifier, also provided as supplementary material (BOP_D_id239_supp_newGT_visualizations.mp4), to convince the reader of their quality.
+These images are taken from a video compilation of all ground truth annotations sorted by object identifier, also provided as supplementary material (BOP_Distrib_id8513_supp_newGT_visualizations.mp4), to convince the reader of their quality.
 We invite the reader to stop on some frames and check that the distribution recovered by our method does correspond to the ambiguities in the image for the object in the bounding box.
-1.2 Differences between Original Annotations and Ours
-Figure 4 visualizes the differences between the poses that are accepted by BOP and the ones accepted when using our annotations.
+2.4 T-LESS [13] Annotation Details
+[…]
+The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object.
+For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 2.5$\times$2.5mm2, $\delta=5mm$ and $\tau_{pix}=30pix$.
+2.5 YCB-V [48] Annotation Details
+[…]
+The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object.
+For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 5.3$\times$5.3mm2.
+The pruning stage was not necessary for YCB-V, as objects are simpler, with much less occlusions.
+2.6 Differences between Original Object-wise BOP Annotations and Our Image-wise Annotations
+Figure 7 for T-LESS [13] and Figure 8 for YCB-V [48] highlight the differences between the poses that are accepted by BOP and the ones accepted when using our annotations.
 For this purpose, for each object, we provide a bar plot.
@@
 The bar plots show the percentages after sorting them, i.e., the bar on the left corresponds to the image with the smallest difference.
-The annotations for some objects remain mostly unchanged.
-But because our method analyses more finely the possible object symmetries, many poses are actually not accepted with our annotations for the other objects.
+Concerning T-LESS, the annotations for some objects remain mostly unchanged.
+But because our method analyses more finely the possible object symmetries, many poses are actually not accepted with our annotations for the other objects (mainly the ’circular’ ones).
+T-LESS is a very good dataset for our per-image pose distribution annotation method as it features objects with complex symmetries as well as a lot inter-object occlusions.
+Concerning YCB-V, some objects such as the mug (object 14 on Figure 8) could yield very interesting symmetries with occlusions [30].
+However, the disambiguating handle of the mug is never occluded.
+Hence the BOP symmetries that tag the mug as unambiguous.
+Overall, YCB-V has less potential for displaying visual ambiguities.
+Most objects are disambiguated by texture and the few ambiguous YCB-V objects do not face sufficient occlusions to become ambiguous (objects 11, 14 and 18 mainly), as depicted by the visualization of the scenes in Figure 9.
+We show here that our per-image annotation method is able to retrieve finer symmetries patterns, which have an effect when evaluating Single Pose Estimation methods as in Table 1.
 Figure 7: Visualization of annotation changes compared to the T-LESS original annotations.
-Top: T-LESS objects with their identifier.
+Top: T-LESS objects with their identifiers.
 Bottom: For each object, we plot the percentages of poses kept from the original annotations by our method over the images (sorted by percentages).
@@
 Only Object 18 is perfectly symmetrical and our method retrieves the same poses as the original annotations.
-For the other objects, in particular the objects with complex symmetry patterns like the first 4 objects.
-Our annotations significantly change the original annotations.
+For the other objects, in particular the objects with complex symmetry patterns like the first 4 objects, our annotations significantly change the original annotations.
+These changes in the annotations explain the score changes for T-LESS in Table 1.
 The objects annotated as ’circular’ in BOP are highlighted in orange.
-Figure 5: Visualizing some BOP-D ground truths as computed by our method.
-Each example features an object of interest, in the bounding box in the left image, and its BOP-D pose distribution, split between the rotation part (center) and translation part (right).
+Figure 8: Visualization of annotation changes compared to the YCB-V original annotations.
+Top: YCB-V objects with their identifiers.
+Bottom: For each object, we plot the percentages of poses kept from the original annotations by our method over the images (sorted by percentages).
+YCB-V assumes full rotational symmetries, while our annotation method captures more complex symmetry patterns.
+Only Object 13 is perfectly symmetrical, both in terms of geometry and texture, and our method retrieves the same poses as the original annotations.
+Object 18 is given as completely symmetrical but our method tags it as non-ambiguous due to its texture.
+Similarly, objects 1, 16, 19, 20 and 21 have few symmetrical poses for BOP annotations where our method keep always only one pose, as the texture disambiguate them.
+These changes in the annotations explain the score changes for YCB-V in Table 1.
+The objects annotated as ’circular’ in BOP are highlighted in orange.
+Figure 9: Visualization of YCB-V scenes.
+The scenes present at most 6 different objects, with one instance of each.
+The objects with potential ambiguities (mainly objects 14 and 18) have limited occlusions.
+YCB-V is not really suited for pose distribution evaluation as no object becomes ambiguous, apart from object 13.
+Figure 10: Visualization of T-LESS scenes.
+The scenes present up to 18 instances of different objects, with a lot of occlusions and with ambiguous objects.
+T-LESS is really suited for pose distribution evaluation.
+(illustration borrowed from [13])
+Figure 11: Visualizing some BOP-Distrib ground truths as computed by our method.
+Each example features an object of interest, in the bounding box in the left image, and its BOP-Distrib pose distribution, split between the rotation part (center) and translation part (right).
 As no object present symmetries in translation here, the translation part of the distribution is always on the same point of $\mathbb{R}^{3}$.
@@
 3.1 T-LESS Pose Estimation and Unseen Objects Pose Estimation
-Recently, the BOP challenge [42] made public the pose estimations results of the methods evaluated in the leaderboard  [^3] https://bop.felk.cvut.cz/leaderboards/.
+Recently, the BOP challenge [43] made public the pose estimation results of the methods evaluated in the leaderboard  [^3] https://bop.felk.cvut.cz/leaderboards/.
 Based on these results and the BOP toolkit  [^4] https://github.com/thodan/bop_toolkit/tree/master in which we implemented the variations of MSSD and MSPD, that use our per-image symmetries patterns instead of BOP global object symmetries, we were able to reprocess these pose estimates against our new ground truths.
-Our results have been presented in Table 1.
+Our results on T-LESS have been presented in Table 1.
 Section 4 of supplementary material illustrates some of the failure cases with the new and more accurate ground truth.
+3.2 YCB-V Pose Estimation
+We conducted a similar experience of reprocessing BOP competitors on our re-annotation of the YCB-V. Our results have been presented in Table 1.
 3.3 Computation of the SpyroPose [9] Distribution Results
+Beyond finer evaluation of Single Pose Estimation methods, our new per-image annotations allow us to propose the first evaluation on real data of Pose Distribution Estimation methods.
+As no pose distribution evaluation existed, the authors of [9] reported only per-object averaged log-likelihood against BOP original ground truth.
 The implementation provided by the authors of SpyroPose allows to train a network for one object of T-LESS.
@@
 As the probabilities of these estimates quickly tend to zero, we reduce their distributions to the 400 best estimates.
-These 400 estimates are used for the evaluations of Section 5.4.
-As no pose distribution evaluation existed, the authors of [9] reported only per-object averaged log-likelihood against BOP original ground truth.
+These 400 estimates are used for the evaluations of Section 5.4 of the article.
 3.4 Computation of the Lie-Pose Diffusion [19] Distribution Results
+Similarly, as no pose distribution evaluation existed, the authors of [19] reported only few qualitative illustrations of pose distribution results on T-LESS.
+They evaluated their method as a Single Pose Estimation method, with a single run of the method.
 The implementation provided by the authors of Lie-Pose allows to train a network for all objects of T-LESS.
@@
 The authors report 1000 runs to produce a distribution.
-For our evaluation, we ran the code 100 times with different noises to produce Lie-Pose distribution results.
-As no pose distribution evaluation existed, the authors of [19] reported only few illustrations of pose distributions results on T-LESS.
-They evaluated their method as a single pose estimation method, with a single run of the method.
+For our evaluation, due to important computation time, we ran the code 100 times with different input noises to produce Lie-Pose distribution results.
 4 Illustrations of Single Pose Re-evaluation
 We provide here more cues about the changes in the Single Pose ranking, based on the MSSD and MSPD metrics using our new more accurate ground truth.
-To do so, we look at the pose estimates of the method GDRNPP v1.4.
-When we evaluated GDRNPP v1.4 estimates against our new ground truths, it changed the method ranking from rank 6 to rank 10 in Table 1.
+To do so, we look at the pose estimates of the method gdrnpp-pbrreal-rgbd-mmodelv1.3.
+When we evaluated gdrnpp-pbrreal-rgbd-mmodelv1.3 estimates against our new ground truths, it changed the method ranking from rank 7 to rank 10 in Table 1.
 Even more interestingly, the metrics went down from an MSSD of 88.4 and an MSPD of 90.9 to an MSSD of 60.6 and an MSPD of 63.7.
 Figure 12 illustrates why MSSD and MSPD changed for this method.
-It appears that, although GDRNPP v1.4 estimates were close to the ground truth pose, its rotations were not precise.
+It appears that, although gdrnpp-pbrreal-rgbd-mmodelv1.3 estimates were close to the ground truth poses, its rotations were not precise.
 When they are evaluated against a symmetries pattern that is not precise, the evaluation appears correct.
-Our new ground truth shows that GDRNPP v1.4 tends not to align correctly some of the objects.
+Our new ground truth shows that gdrnpp-pbrreal-rgbd-mmodelv1.3 tends not to align correctly some of the objects.
 Hence the drop in performances.
 Refer to caption Refer to caption
-(a) BOP continuous symmetry pattern of Object 2 (envelop) and GDRNPP v1.4 estimate (plain circle).
+(a) BOP continuous symmetry pattern of Object 2 (envelop) and gdrnpp-pbrreal-rgbd-mmodelv1.3 estimate (plain circle).
 The circle belongs to the envelop, yielding a low MSSD error of 4.46.
@@
 The circle does not belong to the envelop anymore, the MSSD error becomes 21.62.
-(c) Corresponding image for (a) and (b): T-LESS Scene 1, Image 17, Object 2 (in bounding box).
 Refer to caption Refer to caption
-(d) BOP continuous symmetry pattern for Object 1 (envelop) and GDRNPP v1.4 estimate (plain circle).
+(c) Corresponding image for (a) and (b): T-LESS Scene 1, Image 17, Object 2 (in bounding box) and rendering of the pose (red arrow highlights disambiguating element).
+Refer to caption Refer to caption
+(d) BOP continuous symmetry pattern for Object 1 (envelop) and gdrnpp-pbrreal-rgbd-mmodelv1.3 estimate (plain circle).
 The circle belongs to the envelop, yielding a low MSSD error of 9.43.
@@
 Refer to caption Refer to caption
-BOP continuous symmetry pattern of Object 4 (envelop) and GDRNPP v1.4 estimate (plain circle).
+BOP continuous symmetry pattern of Object 4 (envelop) and gdrnpp-pbrreal-rgbd-mmodelv1.3 estimate (plain circle).
 The circle belongs to the envelop, yielding a low MSSD error of 3.33.
@@
 The circle does not belong to the envelop anymore, the MSSD error becomes 35.38.
-(h) Corresponding image for T-LESS Scene 5, Image 70, Objects 1 (d-e) and and 4 (f-g), in bounding boxes).
+Refer to caption Refer to caption Refer to caption
+(h) Corresponding image for T-LESS Scene 5, Image 70, Objects 1 (d-e) and and 4 (f-g), in bounding boxes) and renderings of the poses (red arrow highlights disambiguating elements).
 Figure 12: Impact of our annotations on Single Pose evaluation.
-We show here cases where the estimates by a state-of-the-art method (GDRNPP v1.4) produces fairly good MSSD errors when considering ground truth provided by BOP.
+We show here cases where the estimates by a state-of-the-art method (gdrnpp-pbrreal-rgbd-mmodelv1.3) produces fairly good MSSD errors when considering ground truth provided by BOP.
 For these cases, our more accurate ground truth yields worse MSSD errors, as it appears that the estimate belongs to the global symmetry pattern, but does not explain what is visible in the image.
 5 Visualizing results by SpyroPose [9]
-We display some of SpyroPose distribution estimates against our ground truth in Figure 7.
+We display some of SpyroPose distribution estimates against our ground truth in Figure 13 on T-LESS.
 For the case of the three instances of Object 1, SpyroPose correctly retrieves the single mode for Instance 1.
 The continuous symmetry of Instance 2 is partially retrieved.
-Figure 7: Illustration of SpyroPose [9] results..
+Figure 13: Illustration of SpyroPose [9] results on T-LESS..
 We show the distribution estimates for the 3 instances of Object 1 (in the bounding box).
@@
 6 Visualizing results by LiePose diffusion [19]
-We show some of LiePose [19] distribution estimates against our ground truth in Figure 8.
+We show some of LiePose [19] distribution estimates against our ground truth in Figure 14 on T-LESS.
 Similarly to SpyroPose [9], LiePose [19] is able to retrieve the single mode of Instance 1, but gets better results when estimating continuous distributions.
-Figures 9, 10, 11 and 12 compare SpyroPose [9] and LiePose [19] results on objects with discrete and continuous symmetries.
+Figures 15, 16, 17 and 18 compare SpyroPose [9] and LiePose [19] results on objects with discrete and continuous symmetries.
 SpyroPose [9] rotations tends to be more precise than LiePose [19], but misses some of the modes.
 LiePose [19] tends to estimate continuous symmetries when the image produces discrete ones.
-These images are taken from a video compilation of SpyroPose [9] and LiePose [19] results also provided as supplementary material (BOP_D_id239_supp_distribution_comparison_SpyroPose_LiePose.mp4).
+These images are taken from a video compilation of SpyroPose [9] and LiePose [19] results also provided as supplementary material (BOP_Distrib_id8513_supp_distribution_comparison_SpyroPose_LiePose.mp4).
 Scenes with single instance of objects have been chosen, to facilitate visualization.
@@
 Our ground truth distribution is displayed as the envelop for the rotation part and as red stars for the translation part.
-Figure 8: Illustration of LiePose [19] results..
+Figure 14: Illustration of LiePose [19] results on T-LESS..
 We show the distribution estimates for the 3 instances of Object 1 (in the bounding box).
@@
 Both methods are able to retrieve the four rotation modes of the object.
-The envelop in the rotation part represents our BOP-D annotation.
+The envelop in the rotation part represents our BOP-Distrib annotation.
 We provide much more examples in the accompanying video.
@@
 Both methods fail to generate the target continuous rotation, although SpyroPose produces more correct rotations.
-The envelop in the rotation part represents our BOP-D annotation.
+The envelop in the rotation part represents our BOP-Distrib annotation.
 We provide much more examples in the accompanying video.
@@
 For this case of six rotations modes, SpyroPose is able to retrieve only two of them, whereas LiePose tends to a continuous distribution, thus generating false rotations.
-The envelop in the rotation part represents our BOP-D annotation.
+The envelop in the rotation part represents our BOP-Distrib annotation.
 We provide much more examples in the accompanying video.
@@
 For this case of eight rotations modes, SpyroPose is able to retrieve only two of them, whereas LiePose tends to a continuous distribution, thus generating false rotations.
-The envelop in the rotation part represents our BOP-D annotation.
+The envelop in the rotation part represents our BOP-Distrib annotation.
 We provide much more examples in the accompanying video.
@@
 We take ITODD’s small validation set, for which the ground truth is public.
-Figure 13 presents our result for the star object new ground truth, as well as one case of the current best performer (gpose2023) failing to align the holes.
+Figure 19 presents our result for the star object new ground truth, as well as one case of the current best performer for Single Pose estimation (gpose2023 [49]) failing to align the holes.
 In the current version of BOP evaluation, this pose estimation is validated.
@@
 Figure 19: Illustration of ITODD symmetries on the validation set (with public GT).
-For the star image (left, first row), BOP symmetries display 12 rotation modes (middle), whereas our method would keep only 2 rotation modes, which align the two holes (right) (size was set to one over the number of modes, hence the bigger modes on the right).
-We also show a pose estimate of GPose [48], ranked first at BOP 2023, for the star object overlayed on its image.
+For the star image (left, first row), BOP symmetries display 12 rotation modes (middle), whereas our annotation method keeps only 2 rotation modes (right), which align the two holes (size was set to one over the number of modes, hence the bigger modes on the right).
+We also show (second row left) a pose estimate of GPose [49], ranked first at BOP 2023, for the star object overlayed on its image.
 We observe that the holes are not correctly aligned.

### v4

@@
 Current annotations in BOP [21] datasets are given as a single pose, shown here as a circle in the SO(3) representations.
-T-LESS [13] also provides a symmetry pattern leaving the 3D object model invariant, from which a distribution can be computed (the colored points in the SO(3) representations).
-Such distribution however does not cover many cases [30]: For the object in this example, when only its core is visible (Case 1), the pose is fully ambiguous and should be represented by a continuous distribution in SO(3).
+BOP also provides a symmetry pattern per object, from which a distribution can be computed (the colored points in SO(3)).
+Such distribution however does not cover many cases [35]: In this example, when only the core is visible (Case 1), the pose is fully ambiguous and should be represented by a continuous distribution in SO(3).
 When the sides of the head are visible (Case 2), there are still ambiguities and the distribution is made of 6 modes.
-When the hole is visible (Case 3), there are no ambiguities, and the pose distribution should be concentrated around one pose.
-Our method annotates the T-LESS dataset with per-image distributions, taking into account the partial occlusions and allowing us to evaluate a predicted pose properly.
-We show that considering these distributions for evaluating state-of-the-art methods results in a significant change of ranking for the BOP challenge.
+When the hole is visible (Case 3), the pose distribution should be concentrated around one non-ambiguous pose.
+Our method annotates scenes with per-image distributions, taking into account the partial occlusions and allowing us to evaluate a predicted pose properly.
+We show that considering these distributions for evaluation results in a significant change of ranking for the BOP challenge.
 Such ground truth distributions also become a key asset when it comes to evaluating pose distribution estimation methods [13, 23].
@@
 grasping or manipulation), augmented reality, industrial quality control, etc.
-In such contexts, datasets of images with ground-truth poses are of primary importance, since they can be used to train learning based methods, but first and foremost to benchmark the performances of proposed approaches.
+In such contexts, datasets of images with ground truth poses are of primary importance, since they can be used to train learning based methods, but first and foremost to benchmark the performances of proposed approaches.
 The annotation accuracy of reference benchmarks, such as BOP [47] for 6D pose estimation, is thus crucial since it influences the research directions.
@@
 Current reference benchmarks [47] ignore these occlusion-induced symmetries, resulting in an imperfect evaluation of the methods.
-As we will show, considering better ground truth distributions results in a significant change of ranking of current pose estimation methods.
-In this paper, we first propose a method that automatically annotates pose distributions for images that take into account the ambiguities due to the object symmetries but also to the partial occlusions.
-Our method exploits the current annotations as a starting point, the ground truth visibility masks of the objects, and their 3D models.
+Considering better ground truth distributions results in a significant change of ranking of pose estimation methods.
+In this paper, we propose a method that automatically annotates pose distributions for images that take into account ambiguities due to object symmetries but also to partial occlusions.
+Our method exploits the current annotations, the ground truth visibility masks of the objects, and their 3D models, to produce a non-parametric distribution representation (see Supp.
+Mat.
+Section 2.2).
 Its genericity is demonstrated on T-LESS [17] and YCB-V [54].
@@
     A new evaluation framework of 6-DOF pose distribution estimation methods, which makes it the first multi-modal pose distribution benchmarking on real data.
-Our annotations and code will be made public.
 2 Related Work
@@
 for a cup [18, 35]).
-While the state of the art mostly focused on estimating a single pose [6, 39, 15, 24, 42, 45, 8, 41], some recent works focus on estimating the complete set of solutions [30, 37, 34, 23, 31, 18, 9, 19].
+While the state of the art mostly focused on estimating a single pose [10, 44, 16, 28, 48, 51, 12, 46], some recent works focus on estimating the complete set of solutions [35, 42, 39, 27, 36, 22, 13, 23, 11, 7, 4, 55, 30, 50].
 As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are then required, even for methods that return a single pose per image.
@@
 The transformation set that relates the different solutions is then specific to each image and cannot be pre-computed from the 3D object model.
-To our knowledge, no method has been proposed to determine this ground-truth transformation set per image.
+To our knowledge, no method has been proposed to determine this ground truth transformation set per image.
 Instead, the current gold standard in 6D pose estimation benchmarks still consists in approximating this per-image transformation set with a unique global transformation set, usually computed with the same method than fully symmetric object with large enough tolerance to ignore small disambiguating elements (e.g.
@@
 Evaluation metrics.
-Performance evaluation usually differs depending on whether the method is used to output a single pose or a distribution.
+Performance evaluation usually differs if the method outputs a single pose or a distribution.
 For single pose estimation, the accuracy is measured through the deviation of the object surface points when transformed by the estimate and by the ground truth pose (registration error) [18].
@@
 robotics or Augmented Reality), deviations can be measured in 3D space (3D distance) or in the image space (reprojection error), and the registration error can be considered as the mean or maximal error over the object surface.
-In case of multi-valued ground-truths due to a symmetrically tagged object, the accuracy corresponds to the minimal error with respect to the set of solutions [12, 48, 44].
+In case of multi-valued ground truths due to a symmetrically tagged object, the accuracy corresponds to the minimal error with respect to the set of solutions [18, 54, 49].
 In practice, the most commonly used accuracy measures are :
@@
   4. 4.
-    Maximum Symmetry-Aware Projection Distance (MSPD) [16] measures the maximum reprojection error between projections of the ground truth and estimated surface points, with symmetry management, and provides an error useful for augmented reality.
+    Maximum Symmetry-Aware Projection Distance (MSPD) [20] measures the maximum reprojection error between projections of the ground truth and estimated surface points, with symmetry management.
 Based on these accuracy measures, and inspired from the evaluation of the detection methods, single 6D pose estimation methods are usually evaluated through Precision/Recall.
@@
 Moreover, whereas single pose and multi-modal distribution pose methods are currently evaluated with non-comparable metrics, we introduce new evaluation metrics for multi-modal methods that are homogeneous with single pose methods (Section 4).
-[…]
+Figure 2: Method overview.
+From a symmetry candidate set, we pre-compute the object per-vertex $\epsilon\text{-sym}$.
+Then for a given scene, we compute the vertices visibility ($\checkmark$ and illustrate respectively if the visibility test passed or not for the vertex) and perform a robust intersection between their $\epsilon\text{-sym}$.
+This intersection is then pruned with a depth comparison and the result constitutes the symmetries pattern of this object instance for this image.
+When multiplied by the ground truth, we obtain the SE(3) distribution of the object instance.
 3 Method
 Consider a 3D object model $M$ observed from a pose $P_{\text{GT}}$.
-We call symmetry-pattern the set ${\cal T}=\{T_{i}\in{\text{SE}}(3)\}$ of rigid transformations that, once combined with $P_{\text{GT}}$, generate representations that are indistinguishable from the one generated by $P_{\text{GT}}$.
+We call symmetry-pattern the set ${\cal T}=\{T_{i}\in{\text{SE}}(3)\}$ of rigid transformations that, once combined with $P_{\text{GT}}$, generate representations that are geometrically indistinguishable from the one generated by $P_{\text{GT}}$.
 Depending on the object geometry $M$ and the representation type, the set $\cal T$} can be limited to the identity transform or can include multiple transformations in addition to the identity.
@@
 where $\epsilon$ is the maximal deviation tolerance, * represents the application of the transformation $T_{i}$ onto the point set ${\cal V}(M)$ and $d$ measures the max deviation between the sets:
-$d(X,M)=\max_{x\in X}||x-\operatorname*{arg\,min}_{m\in M}(||x-m||_{2})||_{2}$.
+[…]
 In this paper, we propose to consider the case of representation corresponding to an image $I$.
@@
 Indeed, since $\epsilon$-symmetries work in 3D space, they do not take into account the potential self or external occlusions that can be induced by the projection process that produces the image.
-Consequently, to estimate ${\cal T}(I)$, our solution compute first ${\cal T}({\cal V}(M))$ then prune the pose visually-inconsistent (section 3.3).
-As illustrated in Figure 2, to achieve this estimation efficiently, our solution decompose the estimation of ${\cal T}({\cal V}(M))$ in a two step process: a first step of pre-computation (section 3.1) independent of ${\cal V}(M)$ and executed only once, and a second step that compute ${\cal T}({\cal V}(M))$ (section 3.2).
+Consequently, to estimate ${\cal T}(I)$, our solution computes ${\cal T}({\cal V}(M))$ then prune the pose visually-inconsistent (Section 3.3).
+As illustrated in Figure 2, to achieve this estimation efficiently, our solution decomposes the estimation of ${\cal T}({\cal V}(M))$ in a two step process: a first pre-computation (Section 3.1) independent of ${\cal V}(M)$ and executed only once, and a second step that computes ${\cal T}({\cal V}(M))$ (Section 3.2).
 3.1 Speeding up the computation of $\epsilon$-sym
@@
 3.3 Refining $\epsilon$-sym: visual occlusions consideration
-[…]
+Since $\epsilon\text{-sym}$ does not consider visual occlusions, ${\cal T}({\cal V}(M))$ may contain a small set of pose that could induce more visible part of $M$ than ${\cal V}(M)$ (see Supp.
+Mat.
+Section 2.1).
+We follow the VSD metric [18] and determine if a pose $T_{i}$ should be pruned by rendering the depth image $D_{T_{i}}$.
+For annotation purposes, the minimal disambiguation element size to reject a candidate pose should be the same for all objects.
+So we remove VSD dependence to object size and count pixels with a depth deviation to ground truth greater than $\delta$ (mm).
+If the number of pixels exceeds $\tau_{pix}$, the pose is pruned.
 $Prune\big(D_{GT},D_{T_{i}},V_{GT},V_{T_{i}},\delta\big)=\\ \sum_{p\in V_{GT}\cup V_{T_{i}}}\begin{cases}0&\text{if}\;p\in V_{GT}\cap V_{T_{i}}\,\wedge\,\big|D_{GT}(p)-D_{T_{i}}(p)\big|<\delta\\ 1&\text{otherwise,}\end{cases}$ (5)
@@
 where $\zeta$ is the color deviation tolerance and with:
-$d_{\text{color}}(X,M)=\max_{x\in X}d_{\text{col}}(x,\operatorname*{arg\,min}_{m\in M}(||x-m||_{2}))$,
+[…]
 with $d_{\text{col}}(x,m)$ the color distance between vertices $x$ and $m$.
@@
 5 Experiments
-In this section, after evaluating the performances of our annotation method (Section 5.2), we present and discuss the impact of our ground-truth annotations and evaluation metrics onto the evaluation of state-of-the-art solutions for both single pose estimation (Section 5.3) and multi-modal pose distribution estimation (Section 5.4).
+In this section, after evaluating the performances of our annotation method (Section 5.2), we present and discuss the impact of our ground truth annotations and evaluation metrics onto the evaluation of state-of-the-art solutions for both single pose estimation (Section 5.3) and multi-modal pose distribution estimation (Section 5.4).
 Since our annotation process and metrics are related to objects with intrinsic or occlusion-induced symmetries, we perform our evaluations on the T-LESS dataset.
-Indeed, among the datasets of the BOP challenge [17], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground-truth annotations: ITODD [7] and HB [22] feature symmetrical objects, with occlusion-induced symmetries, but do not publicly share their ground-truths.
-HOPE [44] and IC-BIN [5] have symmetrical shapes, but texture information disambiguates them completely.
+Indeed, among the datasets of the BOP challenge [21], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground truth annotations: ITODD [9] and HB [26] feature symmetrical objects, with occlusion-induced symmetries, but do not publicly share their ground truths.
+HOPE [49], Omni6D [56] and IC-BIN [8] have symmetrical shapes, but texture information disambiguates them.
 The first experiment focus on a qualitative evaluation of our annotation method (Section 5.2).
@@
 5.1 Implementation details
-[…]
+In our experiments, we sample uniformly surface points from the CAD models with a resolution of 0.5mm.
+For the pre-computation of elementary symmetries patterns for a given object, we use the per-object symmetries pattern given by the BOP challenge [21] as the initial symmetry candidates, with a tolerance factor $\epsilon\text{-sym}$ set to 1mm.
+Following BOP, continuous symmetries are discretized such as the furthest vertex from the axis of symmetry travels not more than 1% of the object diameter.
 Surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object.
-For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating size of roughly 2.5$\times$2.5mm2, $\delta=5mm$ and $\tau_{pix}=30pix$.
-5.2 Qualitative validation of the method
-We illustrate qualitatively the results obtained on some images of the T-LESS dataset in Figure 3.
-We can observe the coherence between the distribution and the visibility of the disambiguating elements.
-Since this dataset does not contain any image with visual ambiguity for the translational part of the pose, the ground truth translation is always reduced to a single 3D point.
-More visualizations are provided in the supplementary material to convince the reader of the quality of our annotations.
+For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating size of roughly $\times2.5$mm2, $\delta=5$mm and $\tau_{pix}=30$pix.
+5.2 Validation of the method
+Figure 3 illustrates the coherence between the distribution and the visibility of the disambiguating elements.
+The ground truth translation is always reduced to a single 3D point as T-LESS exhibit no translational ambiguity.
+Regarding quantitative evaluation of the symmetry pattern accuracy, Table 1 reports the max deviation error (in mm) between the depth image rendered from the GT pose $P_{GT}$ and the ones rendered from the combination of $P_{GT}$ with any pose of the symmetry pattern.
+We observe that the rendered depths are really close.
+$\mu_{\lvert\delta\rvert}$ ($m$) $\text{med}_{\lvert\delta\rvert}$ ($m$) max|δ| (mm) maxVSD maxPrune (pix)
+0.218 5.85e-3 0.497 0.061 30
+Table 1: Quantitative accuracy analysis of our new pose distribution annotations on T-LESS.
+We report the differences between rendered depths at the ground truth pose and all the points of the corresponding distribution.
+$\mu_{\lvert\delta\rvert}$, $\text{med}_{\lvert\delta\rvert}$ and max|δ| are the mean, median and max of the absolute differences, while maxVSD and maxPrune are max errors of VSD@0.05 and of our pruning.
 Figure 3: Visualizations of our ground truth.
-We display SE(3) ground truth distributions for scene 16 of T-LESS [13].
+We display SE(3) ground truth distributions for scene 1 of T-LESS [17].
 Circle on orientation diagram represents the unique ground truth pose provided as input to our method.
-Colors are used to link object instances to their distributions.
-More visualizations are provided in the supplementary material to convince the reader of the quality of our annotations.
+Colors link objects to their distributions.
 BOP [47] (object-wise annotations) BOP-Distrib (our image-wise annotations)
@@
 We also included two methods for the multi-modal pose distribution estimation task [13, 23] for which we kept only the highest confidence mode.
-Since these methods are not part of the BOP challenge, we retrained them.
 Experimental protocol.
@@
 The Loss is explained by inaccurate pose estimates that were mistakenly validated by previous object-wise annotations (see supplementary materials for visualizations).
-It is also interesting to note that the ranking order of variants v1.2, v1.3 and v1.4 of GDRNPP RGB-D (based on different object detectors, training augmentations, and refinement methods) is changed when evaluated with our ground truth.
-It underlines that inaccurate ground truth annotation can induce erroneous conclusions in ablation studies.
 5.4 Pose Distribution Evaluation
@@
 Their distribution is considered as a Dirac distribution on the single estimate.
-Results are given on the complete dataset, and the subset of objects with cylindrical symmetry (1-4, 13-18, 24 and 30, as illustrated in supplementary material).
+Results are given on the complete dataset, and the subset of objects with cylindrical symmetry (1-4, 13-18, 24 and 30).
 Quantitative results are reported on Table 3 whereas qualitative results are provided in the supplementary material.
@@
 6 Limitations
-Our annotation method relies on a geometric analysis of the scene.
-It is possible that the sensor resolution, sensor noise, field of view, or motion blur may also affect the visibility of the disambiguating parts.
+Our annotation method relies on geometric analysis.
+Sensor resolution, sensor noise, field of view or motion blur may also affect disambiguating parts visibility.
 Considering these effects would improve the ground truth even further.
@@
 For simplicity, most of the 6D pose estimation benchmarks rely on a single 6D pose annotation per image, completed by a per-object symmetry pattern to transform this unique pose into a distribution.
-We argued that ignoring the per-image nature of the symmetry pattern is prone to bias in the resulting distribution ground truth and in the resulting performances evaluations.
+We argued that ignoring the per-image nature of the symmetry pattern is prone to bias in the resulting ground truth and performances evaluations.
 We then proposed a method to annotate 6D pose distribution with a per-image analysis of the object symmetries.
@@
 We also introduced metrics to evaluate methods that estimate a pose distribution and provided their first evaluation on real data.
-We believe that these results should encourage the 3D vision community to adopt a new outlook on these methods and how well they work in practice, and should guide future research.
-To facilitate development, we will make our annotations and code publicly available.
+Such pose distribution qualification is crucial for downstream tasks.
+Indeed, access to multiple accurate pose solutions allows the selection of the best one for the task (e.g., considering obstacles) or, when the object is known to be asymmetrical, helps determine if the task can’t be completed from the current viewpoint due to occlusion of the elements that break the symmetry and select the next best view.
+Acknowledgment
+This work was partly funded by the European Union’s Horizon Europe Research and Innovation Program under Grant 101070227 (CONVINCE) and partly funded by the European Union’s Horizon Europe Research and Innovation program under grant agreement nº 101135708 (JARVIS Project).
+The authors would like to thank Rasmus Laurvig Haugaard for providing SpyroPose [13] implementation for baseline evaluation.
 Supplementary Material
@@
 17 return EpsSym, SampledModel
+Listing 1: Offline $\epsilon\text{-sym}$ pre-computation with all vertices visible, pseudo-code of Equation 3 from Section 3.1.
+The geometric ($d_{Geom}$) and colorimetric ($d_{Color}$) distances are implemented as follows:
 […]
-The geometric ($d_{Geom}$) and colorimetric ($d_{Color}$) distances are implemented as follows:
-$d_{Geom}(x,m)=||x-m||_{2}$ and
 $d_{Color}(x,m)<\zeta\Leftrightarrow\min\Big(|h(x)-h(m)|,\\ |h(x)-h(m)-360|,|h(x)-h(m)+360|\Big)<\zeta_{h},\\ \text{AND~}|s(x)-s(m)|<\zeta_{s},\\ \text{AND~}|v(x)-v(m)|<\zeta_{v}.$
@@
 15 return EpsSymImage
-Listing 2: Image annotation, pseudo-code of equation (4) from section 3.2.
+Listing 2: Image annotation, pseudo-code of Equation 4 from Section 3.2.
 1# Input: ModelCAD, EpsSymImage, PoseGT, SensorDepth, $\delta$
@@
 18 return EpsSymImageGlobal
-Listing 3: Image annotation depth post-processing, pseudo-code of equation (4) from section 3.3, based on VSD implementation from BOP toolkit.
+Listing 3: Image annotation depth post-processing, pseudo-code of Equation 4 from Section 3.3, based on VSD implementation from BOP toolkit.
 Section 3.3 does not use the VSD metric, as VSD is normalized by the number of pixels in the union of the visible masks.
@@
 2 Analysis of our BOP-Distrib Annotations
-2.1 Accuracy of the annotations
-Figure 4 illustrate renderings of all poses of a symmetries pattern.
-Figure 4: Variation between poses in a single mode symmetries pattern.
-We display the renderings of the symmetries pattern of object 2 in image 7, scene 1.
-Although the SO(3) visualization may show spread points, these rotations result in similar poses.
 2.1 False visible points and false occluded points detected by the pruning stage
-[…]
-[…]
-[…]
+Figures 4 and 5 present the results of our pruning procedure.
+Each $\epsilon\text{-sym}$ mode is rendered to be compared to the ground truth pose depth rendering.
+The pixel with deviation greater than $\delta$ are counted, and if they too numerous (more than $\tau_{pix}$), the mode is pruned, as in figure 5.
+Figure 4: Depth deviation post-processing analysis. For a given image, we display the depth renderings of the ground truth pose and of one $\epsilon\text{-sym}$ mode (1 here). They align well.
+Figure 5: Depth deviation post-processing analysis.
+For a given image, we display the depth renderings of the ground truth pose and of one $\epsilon\text{-sym}$ mode (12).
+Mode 12 generates several falsely occluded pixels (where the hole should be) and falsely visible pixels (where the hole is but shouldn’t be).
+Mode 12 is rejected by or pruning stage.
+2.2 Choice of Distribution Representation
+Unlike approaches based on Bingham distributions [7, 4, 11], implicit representations [39, 22], Wigner harmonics [30], matrix Fisher distributions [55] or continuous symmetry groups [2], we do not have a continuous distribution representation.
+We represent the ground truth pose distributions in a non-parametric way with a set samples of it.
+Representing a distribution with a large set of samples is common practice and has the advantage of being general: it can represent distributions with no clear analytical form which happen in case of visual ambiguities beyond symmetries, i.e., it can represent distributions with no clear analytical form which happen in case of visual ambiguities beyond symmetries.
+Moreover, a set of samples permits an efficient performance evaluation.
 2.3 Additionnal BOP-Distrib Ground Truths Visualizations
@@
 2.4 T-LESS [17] Annotation Details
-[…]
+In our experiments, we sample surface points from the CAD models with a resolution of 0.5mm.
+For the pre-computation of elementary symmetries patterns for a given object, we use the per-object symmetries pattern given by the BOP challenge [21] as the initial symmetry candidates, with a tolerance factor $\epsilon\text{-sym}$ set to 1mm.
 The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object.
-For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 2.5$\times$2.5mm2, $\delta=5mm$ and $\tau_{pix}=30pix$.
+For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly $\times2.5$mm2, $\delta=5mm$ and $\tau_{pix}=30pix$.
 2.5 YCB-V [54] Annotation Details
-[…]
+In our experiments, we sample surface points from the CAD models with a resolution of 1mm.
+For the pre-computation of elementary symmetries patterns for a given object, we use the per-object symmetries pattern given by the BOP challenge [21] as the initial symmetry candidates, with a tolerance factor $\epsilon\text{-sym}$ set to 2mm.
+The color tolerance is set in the HSV color space to have the chrominance on a single channel (hue) and only luminance on the other two (saturation and value).
+It is empirically set to 4 in hue and 0.1 in saturation and value.
 The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object.
-For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 5.3$\times$5.3mm2.
+For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly $\times5.3$mm2.
 The pruning stage was not necessary for YCB-V, as objects are simpler, with much less occlusions.
@@
 Similarly to SpyroPose [13], LiePose [23] is able to retrieve the single mode of Instance 1, but gets better results when estimating continuous distributions.
-Figures 15, 16, 17 and 18 compare SpyroPose [9] and LiePose [19] results on objects with discrete and continuous symmetries.
+Figures 14, 15, 16 and 17 compare SpyroPose [13] and LiePose [23] results on objects with discrete and continuous symmetries.
 SpyroPose [13] rotations tends to be more precise than LiePose [23], but misses some of the modes.
@@
 We provide much more examples in the accompanying video.
+7 Metrics for evaluation: extended discussion
+Rotation and translation errors are model-independent.
+[18], which led to BOP, states that ”fitness of object surface alignment is the main indicator of object pose quality, model-dependent pose error functions should be therefore preferred.” However, translation and rotation errors, as definer by [18]) with Equation 9 & Equation 10 can be exploited with our definitions of precision and recall over distributions (Equation 7 & Equation 8).
+$\mathbf{d}_{\text{TE}}(\hat{\textbf{t}},\bar{\textbf{t}})=\|\bar{\textbf{t}}-\hat{\textbf{t}}\|_{2},$ (9)
+$\mathbf{d}_{\text{RE}}(\hat{\textbf{R}},\bar{\textbf{R}})=\text{arccos}\left(\text{Tr}(\hat{\textbf{R}}\bar{\textbf{R}}-1)/2\right),$ (10)
+where $\hat{\textbf{R}}$ and $\hat{\textbf{t}}$ are respectively the ground truth rotation and translation, and where $\bar{\textbf{R}}$ and $\bar{\textbf{t}}$ are respectively the estimated rotation and translation.
+Table 4 is a reprocessing of pose distribution estimation method results, with precision and recall over distribution, as defined by Equation 7 & Equation 8, with translation and rotation errors from Equation 9 & Equation 10.
+Methods $\mathbf{P_{RE}}$ $\mathbf{R_{RE}}$ $\mathbf{P_{TE}}$ $\mathbf{R_{TE}}$
+SpyroPose [13] 73.2 69.1 43 66.9
+LiePose [23] 68 91.1 46.2 92.9
+Table 4: Precision/Recall over distribution for separate rotation and translation errors. We reprocess methods pose distribution estimates with decoupled rotation and translation errors.
+8 Downstream Tasks Discussion
+Our work implication on downstream task is two-fold.
+First, as highlighted by the BOP ranking changes in table 2, removing erroneous poses from the BOP ground truth implies a more reliable performance evaluation.
+Indeed, as illustrated by Fig.
+11 of supp.
+mat., performance variations are related to poses that the current BOP ground-truth unduly classify as valid but are properly classified as invalid by our ground truth.
+This is due to overly lax ground truth annotation: some images are annotated as having multiple pose solutions due to the object symmetries whereas disambiguating elements breaking those symmetries are visible in the image.
+For downstream tasks, such as grasping or Augmented Reality, reliable ranking permits to choose the real best performer, and thus a higher success rate of the task.
+The second implication on downstream tasks is related to Table 3, i.e.
+evaluating the ability of a method to determine the complete distribution of poses that explain the observed image.
+For applications such as grasping or Augmented Reality, if the object includes disambiguating elements, it implies that only one pose is valid for the task (e.g.
+a robot that should grab a mug at a specific position on the handle).
+However, if the observed image can be explained by multiple pose, it implies that disambiguating elements are not visible (e.g.
+the handle of the mug is not visible) and the downstream task is impossible to achieve from this unique viewpoint.
+A method that outputs the full distribution of poses provides the downstream task with the ability to determine if the task can be achieved (case of uni-modal distribution) or not (case of multi-modal solution).
+In such situation, a method that outputs a maximum of one pose does not permit to determine if the task can be achieved or not.
+Moreover, in such situation the complete distribution can help to determine the next best viewpoint to make the task feasible.
+Our distribution recall metric 8 evaluates the capacity of the estimation method to retrieve multi-modal distributions.
+This metric is a key indicator of pose distribution estimation methods for downstream tasks usability Figure 18 illustrates the robotic mug handle grabbing case.
+Figure 18: Additional experiment for mug grasping task.
+We display multiple views of mug, associated to our pose distribution annotation.
+We are interested at pose estimation downstream tasks.
+We take the example of robotic grasping.
+We want the robot to grasp the mug by the handle.
+On view 1, the pose distribution is multi-modal, i.e.
+the object pose is ambiguous and the handle position in space cannot be accessed.
+For view 2, 3 and 4, the pose distribution is uni-modal, the image allows to estimate the pose of the mug without ambiguity.
+In such case, the downstream task of robotic handle grasping becomes feasible.
+Now, when it comes to evaluating pose distribution estimation methods against our ground truth, our recall metric evaluates the capacity of the estimation method to retrieve multi-modal distributions.
+This metric is a key indicator of pose distribution estimation methods for downstream tasks usability.
 9 Using Our Per-Image Annotations for Other Pose Estimation Datasets
@@
 With our annotations, it would be penalized.
+New efforts at proposing challenging pose estimation datasets, such as [56], could be processed by our method.
+However, texture disambiguate a lot of the 3D models, and the scenes do not present much occlusion between objects.
 Figure 19: Illustration of ITODD symmetries on the validation set (with public GT).
@@
 MSPD and MSSD metrics validate this estimate, whereas MSPD and MSSD metrics with our new annotations would have penalized it.
-Acknowledgment
-The authors thank Rasmus Laurvig Haugaard for providing SpyroPose [9] implementation for baseline evaluation.
+10 Discussion on Alignist [50]
+Alignist [50] proposes to estimate rotation distribution for ambiguous object shapes from images.
+It was the first method that introduced a solution for supervising the training with a pseudo-ground truth generated rotation distribution.
+To do so, it resorts to a precomputation of such rotation distibution based on ground truth pose, rotation sampling, and SDF (Signed Distance Function) and Surfemb [12] features comparison.
+This precomputation is performed on renderings of single objects, and used to train a double MLP network to infer these distributions.
+The translation part of the pose is not considered.
+The test of the method on T-LESS [17] is conducted following Gilitschenski [11] protocol: it only processes single isolated objects on black background, and is evaluated with log likelihood.
+In contrast, our annotation procedure does not rely on SurfEmb [12] comparisons which results are not guaranteed but it uses geometrical comparisons (see Equation 1).
+Moreover and unlike Alignist [50], our annotation procedure has a rejection mechanism for false visible points and false occluded points, as illustrated in Section 2.1 (see Section 3.3).
+This point is crucial to be able to generate a proper ground truth annotation.
+Finally, our approach does not need to retrain SurfEmb [12] to annotate a new dataset.
+Alignist [50] and other pose distributions estimation methods would benefit from our more accurate pose distributions for their trainings.
+Finally, Alignist [50] method would benefit from our evaluation framework (see Section 5.4).
+For that though, Alignist [50] would need to be tested on the full T-LESS [17] test set (and not just Gilitschenski [11] protocol that excludes external occlusions).
+We could not conduct such tests as the codes and results are not public.